### v1

# Pure C*-algebras

Abstract.
We show that every $C^{*}$-algebra that is $(m,n)$-pure in the sense of Winter is already pure, that is, its Cuntz semigroup is almost unperforated and almost divisible. More generally, we show that even weaker comparison and divisibility properties automatically lead to pureness.
As an application to the non-simple Toms-Winter conjecture, we show that every $C^{*}$-algebra with the Global Glimm Property and finite nuclear dimension is pure.
Key words and phrases:
$C^{*}$-algebras, Cuntz semigroups, comparison, divisibility, Global Glimm Property
2010 Mathematics Subject Classification:
Primary 46L05; Secondary 19K14, 46L80, 46L85.
RA and FP were partially supported by Spanish Research State Agency (grant No.  PID2020-113047GB-I00/AEI/10.13039/501100011033), by the Comissionat per Universitats i Recerca de la Generalitat de Catalunya (grant No. 2021-SGR-01015) and by the Spanish State Research Agency through the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020-001084-M). HT was partially supported by the Knut and Alice Wallenberg Foundation (KAW 2021.0140).
1. Introduction
The groundbreaking classification of amenable factors by Connes and Haagerup [Con76, Haa87] in the 1970s and 80s sparked the Elliott program to classify simple, nuclear $C^{*}$-algebras. A first breakthrough in this direction was obtained with the classification in the purely infinite case by Kirchberg and Phillips [Kir95, Phi00] in the 1990s. In the stably finite case, the game-changing examples of Toms [Tom08] showed that additional regularity properties have to be assumed for $K$-theoretic classification, and building on decades of research by numerous mathematicians, in a series of spectacular results [Win14, EGLN15, TWW17, GLN20], the classification of simple, nuclear $C^{*}$-algebras that are $\mathcal{Z}$-stable and satisfy the universal coefficient theorem (UCT) was accomplished in the 2010s. (See also the recent approach [CGS+23] and the survey [Whi23].)
While the UCT holds in all cases of interest (and possibly it is automatic for all nuclear $C^{*}$-algebras), it is known that $\mathcal{Z}$-stability can fail for simple, nuclear $C^{*}$-algebras, and it becomes highly relevant to determine when precisely this regularity property holds. To investigate this problem, Winter [Win12] introduced the notion of (m,n)-pureness for simple $C^{*}$-algebras, with $m$ and $n$ quantifying certain comparison and divisibility properties among operators, and he showed that every simple $C^{*}$-algebra that has locally finite nuclear dimension (a strong form of nuclearity) and that is $(m,n)$-pure for some $m,n$ is $\mathcal{Z}$-stable. Building on work of Rørdam [Rør04], Winter also showed that every simple, $\mathcal{Z}$-stable $C^{*}$-algebra is $(0,0)$-pure (abbreviated to just pure), which means that it satisfies the strongest of the comparison and divisibility properties.
The concept of $(m,n)$-pureness does not depend on the ideal structure of the algebra, hence may be stated for non-simple $C^{*}$-algebras; see Definition 5.1. However, as we show in this paper, all versions of pureness coincide. Our first main result is (see Theorem 5.8):
Theorem A.
If a $C^{*}$-algebra is $(m,n)$-pure for some $m,n\in{\mathbb{N}}$, then it is pure.
This should be contrasted with the well-known examples of simple, amenable $C^{*}$-algebras that do not have 00-comparison ([Vil98]), or that do not have 00-almost divisibility ([DHTW09]). Thus Theorem A can be interpreted as saying that if a $C^{*}$-algebra is not pure, then it necessarily has poor comparison or poor divisibility properties.
Previously, Theorem A has been obtained for separable, simple, unital $C^{*}$-algebras with locally finite nuclear dimension [Win12], and recently for arbitrary simple $C^{*}$-algebras [APRT24, Theorem 10.5]. Our proof of the general non-simple, non-nuclear case is based on the methods developed in [APRT24] in combination with new techniques about what we term functional divisibility (see Section 4 for details). In turn, these are partly inspired by results in [RT17].
Thus a key ingredient in our approach is the Cuntz semigroup and its technical features (see Section 2 for the precise definition). This is an invariant constructed by Cuntz in [Cun78], which is modelled after the Murray-von Neumann semigroup of projections but using positive elements instead. The version we use is what is commonly called the complete Cuntz semigroup, considered in [CEI08], and technically more suitable as it behaves well with respect to, for example, inductive limits. The power of the Cuntz semigroup was demonstrated in the context of the classification program: A. S. Toms produced in [Tom08] two $C^{*}$-algebras that were indistinguishable by the Elliott invariant and a number of reasonable extensions of it, but could be told apart using their Cuntz semigroup. The salient distinguishing feature was strict comparison of positive elements. Another, more recent use of this semigroup is related to the structure of $C^{*}$-algebras of stable rank one, carried out in [APRT22], where powerful Cuntz semigroup techniques yielded solutions to three structural open problems.
The Toms-Winter conjecture
This prominent conjecture [Win18] aims at establishing a robust class of $C^{*}$-algebras covered by the classification theorem referred to above. It predicts that for a separable, simple, unital, non-elementary, nuclear $C^{*}$-algebra, the following properties are equivalent:
  1. (1)
    $A$ has finite nuclear dimension.
  2. (2)
    $A$ is $\mathcal{Z}$-stable.
  3. (3)
    $A$ has strict comparison of positive elements.
The nuclear dimension was introduced in [WZ10] as a noncommutative generalization of covering dimension. A $C^{*}$-algebra $A$ is said to be $\mathcal{Z}$-stable if it tensorially absorbs the Jiang-Su algebra $\mathcal{Z}$, that is, if $A\cong\mathcal{Z}\otimes A$. A $C^{*}$-algebra $A$ has strict comparison of positive elements if it has 00-comparison; see Definitions 3.1 and 5.1 for details.
In two groundbreaking papers [Win12, CET+21], it was shown that conditions (1) and (2) of the conjecture are equivalent. This was later generalized to stably projectionless, simple $C^{*}$-algebras in [Tik14, CE20], thereby showing that a separable, simple, non-elementary $C^{*}$-algebra has finite nuclear dimension if and only if it is nuclear and $\mathcal{Z}$-stable.
Further, it is known that (2) implies (3) [Rør04]. The implication (3$\Rightarrow$(2) has been verified under certain additional assumptions [Sat12, KR14, TWW15, Thi20], but it remains open in general.
The Toms-Winter conjecture is also closely related to pureness, which we view as a property lying between (2) and (3). Indeed, as mentioned above, $\mathcal{Z}$-stability implies pureness, which in turn implies strict comparison of positive elements; see Proposition 5.2. Further, among separable, simple, unital, non-elementary $C^{*}$-algebras, pureness implies $\mathcal{Z}$-stability under the additional assumption of locally finite nuclear dimension [Win12], and strict comparison of positive elements implies pureness under the additional assumption of a positive solution to the Rank Problem. Such a positive solution has been obtained for $C^{*}$-algebras of stable rank one [Thi20, APRT22].
The non-simple Toms-Winter conjecture
The regularity properties of the Toms-Winter conjecture are also relevant for non-simple $C^{*}$-algebras, and an investigation of their interplay was initiated in [RT17]. While the assumption of unitality in the original formulation of the conjecture could be dispensed with, the condition to be non-elementary is crucial. Indeed, elementary $C^{*}$-algebras have nuclear dimension zero, but they are neither $\mathcal{Z}$-stable nor pure.
There are two natural generalizations of non-elementariness to the non-simple setting: Nowhere scatteredness, and the Global Glimm property. Here, a $C^{*}$-algebra is nowhere scattered if none of its quotients has a nonzero, elementary ideal, or equivalently if every hereditary sub-$C^{*}$-algebra is generated by nilpotent elements [TV24]. A $C^{*}$-algebra has the Global Glimm Property if every hereditary sub-$C^{*}$-algebra contains an approximately full, nilpotent element [TV23]. Clearly, every $C^{*}$-algebra with the Global Glimm Property is nowhere scattered, and the Global Glimm Problem asks whether the converse holds. We ask:
Question B.
Let $A$ be a separable, nowhere scattered, nuclear $C^{*}$-algebra. Are the following properties equivalent?
  1. (1)
    $A$ has finite nuclear dimension.
  2. (2)
    $A$ is $\mathcal{Z}$-stable.
  3. (3a)
    $A$ is pure.
  4. (3b)
    $A$ has strict comparison of positive elements.
The implications (2$\Rightarrow$(3a)$\Rightarrow$(3b) hold in general; see Proposition 5.2. We also note that pureness implies the Global Glimm Property; see [TV23, Theorem 3.6]. Therefore, Question B contains in particular the question of whether the Global Glimm Problem has a positive solution for $C^{*}$-algebras with finite nuclear dimension. As for simple $C^{*}$-algebras, the question whether strict comparison of positive elements implies pureness is closely related to the Rank Problem. This was solved for $C^{*}$-algebras of real rank zero or stable rank one ([ER06] and [APRT22]), which shows that (3b) implies (3a) in these cases.
In [RT17, Theorems 7.10, 7.15] it is shown that (3a) implies (2) under the additional assumption that the $C^{*}$-algebra $A$ has locally finite nuclear dimension, that no quotient of $A$ has a simple, purely infinite ideal, and that the primitive ideal space of $A$ has a basis of compact-open subsets or is Hausdorff.
The following result (which is Theorem 6.4) provides a further partial answer to Question B.
Theorem C.
Every $C^{*}$-algebra with the Global Glimm Property and finite nuclear dimension is pure.
Acknowledgements
The authors thank Leonel Robert for valuable feedback on an earlier version of this paper.
2. Preliminaries
The techniques used in the paper mainly concern the structure of the Cuntz semigroup $\operatorname{Cu}(A)$ of a $C^{*}$-algebra $A$, whose definition and properties we describe below.
Given positive elements $a,b$ in a $C^{*}$-algebra $A$, one says that $a$ is Cuntz subequivalent to $b$, denoted by $a\precsim b$, if there is a sequence $(r_{n})_{n}$ in $A$ such that $\lim_{n\to\infty}\|a-r_{n}br_{n}^{*}\|=0$. Further, $a$ and $b$ are Cuntz equivalent, denoted $a\sim b$, if $a\precsim b$ and $b\precsim a$. These relations were introduced by Cuntz in [Cun78].
[…]
It is proved in [CEI08] that $\operatorname{Cu}(A)$ satisfies the following properties defined for a general positively ordered monoid $S$ (and usually referred to as axioms):
  1. (O1)
    If $(x_{n})_{n}$ is an increasing sequence in $S$, then $\sup_{n}x_{n}$ exists.
  2. (O2)
    For any $x\in S$ there exists a sequence $(x_{n})_{n}$ such that $x_{n}\ll x_{n+1}$ for all $n$ and $x=\sup_{n}x_{n}$. (We say that $(x_{n})_{n}$ is a $\ll$-increasing sequence.)
  3. (O3)
    If $x_{1}\ll x_{2}$ and $y_{1}\ll y_{2}$, then $x_{1}+y_{1}\ll x_{2}+y_{2}$.
  4. (O4)
    If $(x_{n})_{n}$ and $(y_{n})_{n}$ are increasing sequences in $S$, then $\sup_{n}(x_{n}+y_{n})=\sup_{n}x_{n}+\sup_{n}y_{n}$.
The relation $\ll$ in these axioms is defined as follows: $x\ll y$ if for every increasing sequence $(y_{n})_{n}$ satisfying $y\leq\sup_{n}y_{n}$ there exists $n_{0}\in{\mathbb{N}}$ such that $x\leq y_{n_{0}}$. The relation $\ll$ is called the way-below relation, or compact containment relation, and one says that ‘$x$ is way-below $y$’ if $x\ll y$. An element $u\in S$ such that $u\ll u$ is termed compact.
A positively ordered monoid satisfying axioms (O1)-(O4) is called a $\mathrm{Cu}$-semigroup and belongs to a category of semigroups that has thoroughly been studied (see, for example, [APT18], [APT20a], [APT20b] [APT20c], [TV22], [APRT22], among others; see also the recent survey [GP23]). Yet, when dealing with semigroups obtained as Cuntz semigroups of $C^{*}$-algebras, or further restricting to particular families of $C^{*}$-algebras, additional axioms and properties are satisfied.
The first axiom that is added after (O4) is the following:
  1. (O5)
    For all $x^{\prime},x,y$ with $x^{\prime}\ll x\leq y$ there exists $z$ such that $x^{\prime}+z\leq y\leq x+z$. Moreover, if $x+w\leq y$ for some $w$, and $w^{\prime}\ll w$, then $z$ may be chosen such that $w^{\prime}\ll z$.
This axiom is usually referred to as the axiom of almost algebraic order. The first part of the statement above was proved in [RW10, Lemma 7.1], and the second part can be found in [APT18, Proposition 4.6].
Since we will often use the following consequence of (O5), we add a proof here for completeness.
Lemma 2.1 ([TV23, Lemma 2.2]).
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), let $k\in{\mathbb{N}}$, and let $z^{\prime},z,x\in S$ satisfy $z^{\prime}\ll z$ and $(k+1)z\leq x$. Then there exists $d\in S$ such that
$kz^{\prime}+d\leq x\leq kz+d,\,\,\,\text{ and }\,\,\,x\leq(k+1)d.$
Proof.
Choose $z^{\prime\prime}\in S$ such that $z^{\prime}\ll z^{\prime\prime}\ll z$. Applying (O5) for
$kz^{\prime}\ll kz^{\prime\prime},\quad z^{\prime\prime}\ll z,\,\,\,\text{ and }\,\,\,kz^{\prime\prime}+z\leq x,$
we obtain $d\in S$ satisfying
$kz^{\prime}+d\leq x\leq kz^{\prime\prime}+d,\,\,\,\text{ and }\,\,\,z^{\prime\prime}\ll d.$
Then $x\leq(k+1)d$, which shows that $d$ satisfies the desired conditions. ∎
Two more axioms are used very often in the theory of Cu-semigroups. The first one, (O6), is a weaker form of Riesz decomposition and the second one, (O7), is a weak form of interpolation. That the Cuntz semigroup of any $C^{*}$-algebra satisfies both (O6) and (O7) was proved in [Rob13, Proposition 5.1.1] and [APRT21, Proposition 2.2], respectively. We will not need the precise formulation of these axioms here, but rather a very useful consequence of them that allows to take infima with idempotent elements. We make this precise below.
Recall that a $\mathrm{Cu}$-semigroup $S$ is said to be countably based if all elements can be written as suprema of elements from a fixed countable subset. It is known that $\operatorname{Cu}(A)$ is countably based whenever $A$ is separable (for a proof, see [APS11, Lemma 1.3]).
Theorem 2.2 (cf. [APRT21, Theorems 2.4, 2.5]).
Let $S$ be a countably based $\mathrm{Cu}$-semigroup satisfying (O5)-(O7). Then each $x\in S$ and each idempotent element $w\in S$ have an infimum $x\wedge w$ in $S$. Further, the map $S\to S$ given by $x\mapsto x\wedge w$ is a monoid homomorphism preserving the order and the suprema of increasing sequences.
Let $S$ be a $\mathrm{Cu}$-semigroup. A map $\lambda\colon S\to[0,\infty]$ is called a functional if $\lambda$ preserves addition, the zero element, order and suprema of increasing sequences. We denote by $F(S)$ the set of functionals on $S$. This is a cone when endowed with the operations of pointwise addition and pointwise scalar multiplication by positive real numbers. When equipped with the appropriate topology, $F(S)$ becomes a compact Hausdorff cone (see [ERS11, Theorems 4.4, 4.8], [Rob13, Section 2.2]; see also [Kei17, Theorem 3.17]).
Given $x\in S$, we obtain a map $\widehat{x}\colon F(S)\to[0,\infty]$ defined by $\widehat{x}(\lambda)=\lambda(x)$, for $\lambda\in F(S)$. $\widehat{x}$ is linear and lower semicontiuous, and if $S$ satisfies (O5) and (O6), then $\textrm{Lsc}(F(S))$, the set of all such maps, is an algebraically ordered, lattice ordered, compact cone.
Two main questions arise in this setting. To determine the range of the so-called rank map $x\mapsto\widehat{x}$ (referred to as the Rank Problem); and determine to what extent the order in $S$ can be inferred by the order in $\textrm{Lsc}(F(S))$. Note that $x\mapsto\widehat{x}$ is additive and preserves sequential suprema; it does not preserve the relation of compact containment, and is certainly not an order embedding. However, the result below will be useful.
Given elements $x$ and $y$ in a partially ordered semigroup, one writes $x<_{s}y$ if there exists $k\in{\mathbb{N}}$ such that $(k+1)x\leq ky$. Item (ii) in the following lemma is a consequence of [Rob13, Proposition 2.2.2] or [APT18, Theorem 5.2.19].
Proposition 2.3 ([Rob13, Lemma 2.2.5], [APRT24, Lemma 5.3]).
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), and let $x,y\in S$. Then:
  1. (i)
    If $x\ll y$, and $s,t\in(0,\infty]$ satisfy $s<t$, we have $s\widehat{x}\ll t\widehat{y}$ in $\operatorname{Lsc}(F(S))$.
  2. (ii)
    If $\widehat{x}\leq\gamma\widehat{y}$ for some $\gamma\in(0,1)$, then $x^{\prime}<_{s}y$ for all $x^{\prime}\ll x$.
3. Variants of comparison
We first recall the definition of $m$-comparison as considered by Toms and Winter in [TW09, Lemma 6.1], and later formalized by Ortega, Perera and Rørdam in [OPR12, Definition 2.8].
Definition 3.1.
Let $m\in{\mathbb{N}}$. A $\mathrm{Cu}$-semigroup $S$ has $m$-comparison if for all $x,y_{0},\ldots,y_{m}\in S$ satisfying $x<_{s}y_{j}$ for $j=0,\ldots,m$, we have $x\leq y_{0}+\ldots+y_{m}$. The lowest possible value for $m$-comparison is attained when $m=0$, which is commonly referred to as saying that $S$ is almost unperforated.
We say that a $C^{*}$-algebra has $m$-comparison if its Cuntz semigroup does.
Remarks 3.2.
(1) L. Robert showed in [Rob11, Theorem 1.3] that, if $A$ is a $C^{*}$-algebra with nuclear dimension at most $m$, then $A$ has $m$-comparison.
(2) Let $m\in{\mathbb{N}}$ and let $S$ be a $\mathrm{Cu}$-semigroup that has $m$-comparison. Then, $x<_{s}y$ implies $x\leq(m+1)y$, for all $x,y\in S$.
Given $N\in{\mathbb{N}}$ and elements $x$ and $y$ in a partially ordered semigroup, we write $x\leq_{N}y$ if $nx\leq ny$ for all $n\in{\mathbb{N}}$ with $n\geq N$; see [APRT24, Notation 5.2] and also [APT18, Definition 5.6.1]. To recall the definition of locally bounded comparison amplitude, we need the notion of a scale in a $\mathrm{Cu}$-semigroup.
3.3Scales.
Let $S$ be a $\mathrm{Cu}$-semigroup. A scale in $S$ is a subset $\Sigma\subseteq S$ such that
  1. (i)
    is downward hereditary: if $x\leq y$ for some $x\in S$ and $y\in\Sigma$, then $x\in\Sigma$,
  2. (ii)
    is closed under suprema of increasing sequences, and
  3. (iii)
    $\Sigma$ generates $S$ as an ideal, that is, for $x^{\prime},x\in S$ with $x^{\prime}\ll x$, there exist $y_{1},\ldots,y_{n}\in\Sigma$ such that $x^{\prime}\leq y_{1}+\ldots+y_{n}$.
(See [APT20c, Definition 4.1]). We say that the pair $(S,\Sigma)$ is a scaled $\mathrm{Cu}$-semigroup.
Given a scale $\Sigma$ and $d\in{\mathbb{N}}$, the $d$-fold amplification of $\Sigma$ is defined as $\Sigma^{(0)}:=\{0\}$, and
[…]
for $d\geq 1$.
If $A$ is a $C^{*}$-algebra, then
$\Sigma_{A}:=\big{\{}x\in\operatorname{Cu}(A):x\leq[a]\text{ for some $a\in A_{+}$}\big{\}}$
is a scale in $\operatorname{Cu}(A)$, and the pair $(\operatorname{Cu}(A),\Sigma_{A})$ is called the scaled Cuntz semigroup of $A$; see [APT20c, Paragraph 4.2] and [TV23, Lemma 3.3], and also [APRT24, Paragraph 4.5]. By [APRT24, Proposition 7.4], for $d\geq 1$ we have
$\Sigma_{A}^{(d)}=\big{\{}x\in\operatorname{Cu}(A):x\leq[a]\text{ for some $a\in M_{d}(A)_{+}$}\big{\}}.$
We recall below the definition of locally bounded comparison amplitude, from [APRT24]. Useful reformulations of this condition will be given in Lemma 3.5.
Definition 3.4 ([APRT24, Definition 6.1]).
A scaled $\mathrm{Cu}$-semigroup $(S,\Sigma)$ has locally bounded comparison amplitude, in short (LBCA), if for every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $N=N(\gamma,d)$ such that the following holds:
$\widehat{x}\leq\gamma\widehat{y}\quad\textrm{ implies }\quad x\leq_{N}y,\quad\text{ for all }x,y\in\Sigma^{(d)}.$
A $C^{*}$-algebra will be said to have locally bounded comparison amplitude if $(\operatorname{Cu}(A),\Sigma_{A})$ has (LBCA).
Lemma 3.5.
Let $(S,\Sigma)$ be a scaled $\mathrm{Cu}$-semigroup. The following are equivalent:
  1. (i)
    $(S,\Sigma)$ has locally bounded comparison amplitude.
  2. (ii)
    For every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $N=N(\gamma,d)$ such that
    $\widehat{x}\leq\gamma\widehat{y^{\prime}},\quad y^{\prime}\ll y\quad\text{ implies }\quad Nx\leq Ny,\qquad\text{for all }x,y^{\prime},y\in\Sigma^{(d)}.$
  3. (iii)
    For every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $M=M(\gamma,d)$ such that:
    $\widehat{x}\leq\gamma\widehat{y}\quad\text{ implies }\quad(M+1)x\leq My,\qquad\text{for all }x,y\in\Sigma^{(d)}.$
Proof.
It is clear that (i) implies (ii).
Let us show that (ii) implies (iii). Fix $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$. Choose $\gamma^{\prime}$ and $n\in{\mathbb{N}}$ such that
$\gamma<\gamma^{\prime},\quad\frac{n+1}{n}\gamma^{\prime}<1.$
Let $N=N(\tfrac{n+1}{n}\gamma^{\prime},d(n+1))$ as provided by (ii). We claim that $M=Nn$ has the desired properties to verify (iii) (for $\gamma$ and $d$).
To this end, let $x,y\in\Sigma^{(d)}$ satisfy $\widehat{x}\leq\gamma\widehat{y}$. Let $x^{\prime}\in S$ be such that $x^{\prime}\ll x$. By Proposition 2.3 (i) applied to $x^{\prime}\ll x$ and $1<\frac{\gamma^{\prime}}{\gamma}$, we have that $\widehat{x^{\prime}}\ll\gamma^{\prime}\widehat{y}$. Hence, there exists $y^{\prime}\ll y$ such that $\widehat{x^{\prime}}\leq\gamma^{\prime}\widehat{y^{\prime}}$. Then
$(n+1)\widehat{x^{\prime}}\leq(n+1)\gamma^{\prime}\widehat{y^{\prime}}=\left(\frac{n+1}{n}\gamma^{\prime}\right)n\widehat{y^{\prime}}.$
Since $(n+1)x^{\prime},ny^{\prime}$ and $ny$ belong to $\Sigma^{((n+1)d)}$, it follows from the choice of $N$ that $N(n+1)x^{\prime}\leq Nny$. Passing to the supremum over all $x^{\prime}\ll x$ we get that $N(n+1)x\leq Nny$, and thus $(M+1)x\leq My$.
Finally, note that if elements $x$ and $y$ in a partially ordered semigroup satisfy $(M+1)x\leq My$ for some $M\in{\mathbb{N}}$, then there exists $N\in{\mathbb{N}}$ such that $(n+1)x\leq ny$ for all $n\geq N$, and in particular $x\leq_{N}y$; see [APT18, Proposition 5.2.13]. This shows that (iii) implies (i). ∎
The following lemma is implied by [APRT24, Proposition 10.3].
Lemma 3.6.
Any $C^{*}$-algebra with $m$-comparison has locally bounded comparison amplitude.
Proof.
In order to apply [APRT24, Proposition 10.3] it is required that the Cuntz semigroup of a $C^{*}$-algebra satisfies (O5), (O6), and the so-called Edward’s condition. The first two are always satisfied, as mentioned in the preliminaries. As for the third condition, this was also shown to hold in general in [APRT21, Theorem 5.3]. ∎
Theorem 3.7.
If a $C^{*}$-algebra has finite nuclear dimension, then its Cuntz semigroup has locally bounded comparison amplitude.
Proof.
Let $A$ be a $C^{*}$-algebra with nuclear dimension at most $m$ for some $m\in{\mathbb{N}}$. As observed in Remarks 3.2, $\operatorname{Cu}(A)$ has $m$-comparison, hence the result follows from Lemma 3.6. ∎
4. Variants of divisibility
We first recall the definition of $n$-almost divisibility for $\mathrm{Cu}$-semigroups, which appeared in [RT17, Section 2.3]. Other related notions of divisibility will be also useful in our analysis. These have sprung into prominence due to their connection with important structural properties of $C^{*}$-algebras.
Definition 4.1.
Let $k\in{\mathbb{N}}$, and $n\in{\mathbb{N}}\cup\{\infty\}$. We say that an element $x$ in a $\mathrm{Cu}$-semigroup $S$ is $(k,n)$-divisible if, for each $x^{\prime}\in S$ with $x^{\prime}\ll x$, there exists $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)(n+1)y.$
If all elements are $(k,n)$-divisible, when $n=\infty$, then we say that $S$ is $(k,\omega)$-divisible.
If all elements are $(k,n)$-divisible for all $k$, we then say that $S$ is $n$-almost divisible. In case $S$ is 00-almost divisible, then we just say that $S$ is almost divisible.
A $C^{*}$-algebra is $n$-almost divisible if its Cuntz semigroup is.
Remarks 4.2.
(1) It was proved in [TV24, Theorem 9.1] that, if $A$ is a nowhere scattered $C^{*}$-algebra of real rank zero, then $A$ is almost divisible. In fact, its Cuntz semigroup $\operatorname{Cu}(A)$ is weakly divisible: for any $x\in\operatorname{Cu}(A)$, there are $y,z\in\operatorname{Cu}(A)$ such that $x=2y+3z$; see also [APGPSM10, Section 6]. In the simple case, and for the Murray-von Neumann semigroup, this was partly obtained in [PR04, Proposition 5.3].
(2) If a $\mathrm{Cu}$-semigroup $S$ is almost unperforated, then it is almost divisible precisely when, given $x\in S$ and $k$, there is $y\in S$ such that $ky\leq x\leq(k+1)y$. This was proved in [APT18, Proposition 7.3.7].
(3) The concept of $(k,\omega)$-divisibility is connected to the Global Glimm Property which, in the context of this paper, will be analysed with more detail in Section 6. It was proved in [TV23, Theorem 3.6] that a $C^{*}$-algebra has the Global Glimm Property precisely when its Cuntz semigroup is $(2,\omega)$-divisible.
Inspired by the results in [RT17, Section 6], we introduce the following more general divisibility conditions in terms of comparison by functionals.
Definition 4.3.
We say that an element $x$ in a $\mathrm{Cu}$-semigroup $S$ is
  • (i)
    functionally almost divisible, or that $x$ has (FAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, every $k\in{\mathbb{N}}_{>0}$, and every $\varepsilon>0$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq k\widehat{y}+\varepsilon\widehat{x}.$
  • (ii)
    lower functionally almost divisible, or that $x$ has (LFAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+1\right)\widehat{y}.$
  • (iii)
    upper functionally almost divisible, or that $x$ has (UFAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ and $\nu>0$ such that
    $y\leq x,\quad(k+\nu)\widehat{y}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)y.$
A $\mathrm{Cu}$-semigroup $S$ is said to be (upper, lower) functionally almost divisible if every element in $S$ is (upper, lower) functionally almost divisible.
We begin our study of functional divisibility by providing a number of technical results relating each of the notions above. These results will be applied to $C^{*}$-algebras with $n$-almost divisibility at the end of the section, where we will show that any $n$-almost divisible $C^{*}$-algebra satisfies all three divisibility notions; see Theorem 4.7.
Lemma 4.4.
Let $S$ be a $\mathrm{Cu}$-semigroup.
  1. (i)
    $S$ is functionally almost divisible if and only if for every $x^{\prime},x\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq k\widehat{y}+\tfrac{1}{k}\widehat{x}.$
  2. (ii)
    $S$ is lower functionally almost divisible if and only if for every $x^{\prime},x\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq(k+\tfrac{1}{k})\widehat{y}.$
Proof.
We will only prove (i) since the proof of (ii) is similar. Moreover the forward implications are clear. To prove the backwards implication, let $x^{\prime},x\in S$ satisfy $x^{\prime}\ll x$, let $k\in{\mathbb{N}}_{>0}$, and let $\varepsilon>0$. Choose $n\in{\mathbb{N}}_{>0}$ such that $\tfrac{1}{kn}\leq\varepsilon$. Applying the assumption for $x^{\prime},x$ and $kn$, we obtain $z\in S$ such that
$knz\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq kn\widehat{z}+\frac{1}{kn}\widehat{x}.$
Set $y:=nz$. Then
$ky=knz\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq kn\widehat{z}+\frac{1}{kn}\widehat{x}=k\widehat{y}+\frac{1}{kn}\widehat{x}\leq k\widehat{y}+\varepsilon\widehat{x},$
as desired. ∎
Lemma 4.5.
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5) and (O7). Assume that $S$ is $(2,\omega)$-divisible and is functionally almost divisible. Then $S$ is lower functionally almost divisible.
Proof.
Let $k\in{\mathbb{N}}$, and let $x^{\prime},x\in S$ be such that $x^{\prime}\ll x$. It suffices to prove that there exists $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+\frac{1}{k}\right)\widehat{y}.$
For $k=1$ use $y:=x$. Thus, we may from now on assume that $k\geq 2$.
Using the methods from [TV21, Section 5], we find a countably based, $(2,\omega)$-divisible $\mathrm{Cu}$-semigroup $H$ that contains $x^{\prime},x$ and such that $H$ satisfies (O5), (O7), and (FAD).
Choose $x^{\prime\prime}\in H$ and $n\in{\mathbb{N}}_{>0}$ such that
$x^{\prime}\ll x^{\prime\prime}\ll x,\quad k+1\leq n,\,\,\,\text{ and }\,\,\,\frac{k(k+1)+n}{n}\leq 1+\frac{1}{k^{2}}.$
Since $H$ is $(2,\omega)$-divisible, it is also $(n,\omega)$-divisible by the arguments in [TV23, Lemma 3.4]. Applied to $x^{\prime\prime}\ll x$ and $n$, we obtain $c\in H$ such that
$nc\leq x,\,\,\,\text{ and }\,\,\,x^{\prime\prime}\ll\infty c.$
Using that $x^{\prime\prime}\ll\infty c$, we can choose $c^{\prime}\in H$ such that
$x^{\prime\prime}\ll\infty c^{\prime},\,\,\,\text{ and }\,\,\,c^{\prime}\ll c.$
We have $(k+1)c\leq nc\leq x$. Using Lemma 2.1, we obtain $d\in H$ such that
$kc^{\prime}+d\leq x\leq kc+d,\,\,\,\text{ and }\,\,\,x\leq(k+1)d.$
Thus, one has
$nx^{\prime\prime}\leq nx\leq nkc+nd\leq kx+nd\leq(k(k+1)+n)d.$
Since $H$ is countably based and satisfies (O7), we may apply Theorem 2.2. Hence, for every $s\in S$ the infimum $s\wedge\infty c^{\prime}$ exists, and the map $s\mapsto s\wedge\infty c^{\prime}$ is additive; Set $e:=d\wedge\infty c^{\prime}$. Since $x^{\prime\prime}\ll\infty c^{\prime}$, we have
$nx^{\prime}\ll nx^{\prime\prime}=(nx^{\prime\prime})\wedge(\infty c^{\prime})$ $\leq(k(k+1)+n)d\wedge(\infty c^{\prime})$
$=\big{(}k(k+1)+n\big{)}\big{(}d\wedge(\infty c^{\prime})\big{)}=(k(k+1)+n)e.$
This allows us to find $e^{\prime}\in H$ satisfying
$nx^{\prime}\leq(k(k+1)+n)e^{\prime},\,\,\,\text{ and }\,\,\,e^{\prime}\ll e.$
Let us further choose $e^{\prime\prime}$ such that $e^{\prime}\ll e^{\prime\prime}\ll e$.
There exists $N\in{\mathbb{N}}$ such that $e^{\prime\prime}\leq Nc^{\prime}$. Choose $\varepsilon>0$ small enough such that
$\left(1+\frac{1}{k^{2}}\right)\varepsilon\leq\frac{1}{N}.$
Applying (FAD) to $e^{\prime}\ll e^{\prime\prime}$, $k$ and $\varepsilon$, we obtain $f\in H$ such that
$kf\leq e^{\prime\prime},\,\,\,\text{ and }\,\,\,\widehat{e^{\prime}}\leq\widehat{kf}+\varepsilon\widehat{e^{\prime\prime}}.$
Set $y:=c^{\prime}+f$. We have $kf\leq e\leq d$ and therefore
$ky=kc^{\prime}+kf\leq kc^{\prime}+d\leq x.$
To verify that $\widehat{x^{\prime}}\leq(k+1/k)\widehat{y}$, let $\lambda\in F(H)$. Then
$\begin{split}\lambda(x^{\prime})&\leq\frac{k(k+1)+n}{n}\lambda(e^{\prime})\leq\left(1+\frac{1}{k^{2}}\right)\lambda(e^{\prime})\\ &\leq\left(1+\frac{1}{k^{2}}\right)\left(k\lambda(f)+\varepsilon\lambda(e^{\prime\prime})\right)\leq\left(k+\frac{1}{k}\right)\lambda(f)+\frac{1}{N}\lambda(e^{\prime\prime})\\ &\leq\left(k+\frac{1}{k}\right)\lambda(f)+\lambda(c^{\prime})\leq\left(k+\frac{1}{k}\right)\lambda(y),\end{split}$
as desired. ∎
Lemma 4.6.
Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), and let $x\in S$ be a compact element. If $x$ is functionally almost divisible, then for every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
$y\leq x,\quad(k+\tfrac{1}{2})\widehat{y}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x\leq(k+1)y.$
In particular, $x$ is upper functionally almost divisible.
Proof.
Let $k\in{\mathbb{N}}$. Using (FAD) with $k+1$ and $\varepsilon=1/3(2k+1)(k+1)$, we obtain an element $z\in S$ such that
$(k+1)z\leq x,\,\,\,\text{ and }\,\,\,\widehat{x}\leq(k+1)\widehat{z}+\frac{1}{3(2k+1)(k+1)}\widehat{x}.$
Find $\delta>0$ such that
$\frac{k+1}{1-\delta}\leq(k+1)+\frac{1}{2(2k+1)},\,\,\,\text{ and }\,\,\,\frac{\varepsilon}{(1-\delta)}\leq\frac{1}{2(2k+1)(k+1)}.$
Now, using that $x\ll x$, it follows from Proposition 2.3 (i) that
$(1-\delta)\widehat{x}\ll\widehat{x}\leq(k+1)\widehat{z}+\varepsilon\widehat{x}.$
This allows us to choose $z^{\prime}\ll z$ such that
(1) $\widehat{x}\leq\frac{k+1}{(1-\delta)}\widehat{z^{\prime}}+\frac{\varepsilon}{(1-\delta)}\widehat{x}\leq\left((k+1)+\frac{1}{2(2k+1)}\right)\widehat{z^{\prime}}+\frac{1}{2(2k+1)(k+1)}\widehat{x}$
Using Lemma 2.1, we can now find $y\in S$ such that
$kz^{\prime}+y\leq x\leq(k+1)y.$
Take $\lambda\in F(S)$. If $\lambda(z^{\prime})=\infty$, we have $\lambda(y)=\lambda(x)=\infty$ and so, in particular, $(k+\tfrac{1}{2})\lambda(y)\leq\lambda(x)$.
Thus, assume that $\lambda(z^{\prime})<\infty$. Since $k\lambda(y)\leq k(\lambda(x)-k\lambda(z^{\prime}))$, using (1) at the first step, and $(k+1)z^{\prime}\leq x$ at the last step, one has:
$\begin{split}(k+\tfrac{1}{2})\lambda(y)&\leq(k+\tfrac{1}{2})\left(\left((k+1)+\tfrac{1}{2(2k+1)}\right)\lambda(z^{\prime})+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)-k\lambda(z^{\prime})\right)\\ &=(k+\tfrac{1}{2})\left(\left(1+\tfrac{1}{2(2k+1)}\right)\lambda(z^{\prime})+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)\right)\\ &\leq(k+\tfrac{1}{2})\left(\tfrac{1}{k+1}\left(1+\tfrac{1}{2(2k+1)}\right)\lambda(x)+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)\right)=\lambda(x),\end{split}$
as desired.
The last assertion is clear. ∎
Theorem 4.7.
Let $A$ be a $C^{*}$-algebra. Assume that $A$ is $n$-almost divisible for some $n\in{\mathbb{N}}$. Then $\operatorname{Cu}(A)$ is (upper, lower) functionally almost divisible.
Proof.
Assume that $A$ is $n$-almost divisible. Without loss of generality, we may assume $A$ to be stable. Further, note that we only need to prove that $\operatorname{Cu}(A)$ has (FAD) and (UFAD). Indeed, any $n$-almost divisible $C^{*}$-algebra has a $(2,\omega)$-divisible Cuntz semigroup. Thus, it follows from Lemma 4.5 that $\operatorname{Cu}(A)$ has (LFAD) whenever it has (FAD).
Let $\mathcal{U}$ be a free ultrafilter on ${\mathbb{N}}$, and let $a\in A_{+}$. Set $C=\{a\}^{\prime}\cap A_{\mathcal{U}}$ and $I=\{a\}^{\perp}\cap A_{\mathcal{U}}$. Then, it follows from [RT17, Theorem 6.1] that the class of the unit $[1]$ in $\operatorname{Cu}(C/I)$ has (FAD). Therefore, $[1]$ has (UFAD) by Lemma 4.6.
Now choose $k\in{\mathbb{N}}_{>0}$. Applying (FAD) to $[1]\ll[1]$ with $\varepsilon=\tfrac{1}{k+1}$, we obtain an element $e\in C/I$ such that
$k[e]\leq[1],\,\,\,\text{ and }\,\,\,\widehat{[1]}\leq k\widehat{[e]}+\tfrac{1}{k+1}\widehat{[1]}.$
Setting $\delta=\tfrac{1}{k(k+1)}$, we have $(1+\delta)\widehat{k[1]}\leq k^{2}\widehat{[e]}+\widehat{[1]}$. Now we apply Proposition 2.3(ii) to find a positive integer $s$ such that $(s+1)k[1]\leq s(k^{2}[e]+[1])$.
Separately, we apply (UFAD) to $[1]\ll[1]$ and we obtain an element $\tilde{f}\in C/I$ such that
$[\tilde{f}]\leq[1],\quad(k+\nu)\widehat{[\tilde{f}]}\leq\widehat{[1]},\,\,\,\text{ and }\,\,\,[1]\leq(k+1)[\tilde{f}],$
for some $10<\nu<$ 1. Since $[1]\ll[1]$, we can find $f\in C/I$ with $[f]\ll[\tilde{f}]$ such that $[1]\leq(k+1)[f]$. Again, applying Proposition 2.3(ii), we find a positive integer $t$ such that $(t+1)k[f]\leq t[1]$.
Let $\bar{e},\bar{f}\in C$ be lifts of $e,f$ respectively. Then, there exists $z\in\operatorname{Cu}(I)$ such that
$k[\bar{e}]\leq[1]+z,\quad(s+1)k[1]\leq s(k^{2}[\bar{e}]+[1])+z,$
and
$[\bar{f}]\leq[1]+z,\quad(t+1)k[\bar{f}]\leq t[1]+z,\quad[1]\leq(k+1)[\bar{f}]+z$
in $\operatorname{Cu}(C)$.
Set $b=a\bar{e}$ and $c=a\bar{f}$ in $A_{\mathcal{U}}$. Using that every element in $C$ commutes with $a$ (in particular, those elements applying the Cuntz subequivalences above) and that every element in $I$ is orthogonal to $a$, one obtains
$k[b]\leq[a],\quad(s+1)k[a]\leq s(k^{2}[b]+[a]),$
and
$[c]\leq[a],\quad(t+1)k[c]\leq t[a],\quad[a]\leq(k+1)[c]$
in $\operatorname{Cu}(A_{\mathcal{U}})$.
This shows that $[a]\in\operatorname{Cu}(A_{\mathcal{U}})$ has (FAD) and (UFAD). Now, using the fact that functional comparison can be restated in terms of comparison in the semigroup ([Rob13, Proposition 2.2.2]), a standard model theoretic argument shows that $[a]\in\operatorname{Cu}(A)$ also has (FAD) and (UFAD); see [FHL+21, Proposition 8.1.6] and also [APRT24, Theorem 10.5]. ∎
5. Pure C*-algebras
In this section, we show that every $(m,n)$-pure $C^{*}$-algebra is pure; see Theorem 5.8.
Definition 5.1.
Given $m,n\in{\mathbb{N}}$, we say that a $\mathrm{Cu}$-semigroup $S$ is $(m,n)$-pure if $S$ has $m$-comparison (Definition 3.1) and is $n$-almost divisible (Definition 4.1). We say that $S$ is pure if it is $(0,0)$-pure.
A $\mathrm{Cu}$-semigroup $S$ has strict comparison if, whenever $x,y\in S$ with $x\leq\infty y$ satisfy that $\lambda(x)<\lambda(y)$ for any $\lambda\in F(S)$ normalized at $y$, one has $x\leq y$.
A $C^{*}$-algebra is $(m,n)$-pure if its Cuntz semigroup is. Similarly, a $C^{*}$-algebra is pure if it is $(0,0)$-pure. We also say that a $C^{*}$-algebra has strict comparison of positive elements (by $2$-quasitraces) provided its Cuntz semigroup has strict comparison.
It was proved in [ERS11, Proposition 6.2] that $S$ has strict comparison if and only if it is almost unperforated.
Proposition 5.2.
Every $\mathcal{Z}$-stable $C^{*}$-algebra is pure. Every pure $C^{*}$-algebra has strict comparison of positive elements.
Proof.
Let $A$ be a $\mathcal{Z}$-stable $C^{*}$-algebra. It was shown in [Rør04, Theorem 4.5] that $\operatorname{Cu}(A)$ is almost unperforated, and thus by the comment above $A$ has strict comparison of positive elements. The argument used to prove almost unperforation can be used to show almost divisibility. For a proof, see e.g. [APT11, Theorem 5.35]. ∎
Remark 5.3.
It is perhaps worth pointing out that the results in [Rør04] deal with the uncomplete Cuntz semigroup $\mathrm{W}(A)$. This makes no difference regarding almost unperforation, since $\operatorname{Cu}(A)$ is a completion of $\mathrm{W}(A)$ (see [APT18, Theorem 3.2.8]), hence $\mathrm{W}(A)$ is almost unperforated if and only if $\operatorname{Cu}(A)$ is.
One can ask whether, for $\mathrm{Cu}$-semigroups, $(0,0)$-pureness agrees with some sort of tensorial absorption. This was explored in [APT18], together with an extensive analysis of the tensor product in the category $\operatorname{Cu}$.
Theorem 5.4 ([APT18, Theorem 7.3.11]).
A $\mathrm{Cu}$-semigroup $S$ is pure if and only if $S\cong\operatorname{Cu}(\mathcal{Z})\otimes S$.
If the Cuntz semigroup of a $C^{*}$-algebra $A$ is isomorphic to that of $\mathcal{Z}\otimes A$, then $A$ is clearly pure. In the proof of [Tom11, Theorem 1.2], Toms shows that the converse holds for simple $C^{*}$-algebras. (See also [APP18, Section 7].) The following question is thus pertinent:
Question 5.5.
Let $A$ be a pure $C^{*}$-algebra. Is $\operatorname{Cu}(A)\cong\operatorname{Cu}(\mathcal{Z}\otimes A)$?
We start with an important technical result, which generalizes some of the $\mathrm{Cu}$-semigroup techniques underlying Proposition 6.4 and Theorem 10.5 in [APRT24].
Proposition 5.6.
Let $(S,\Sigma)$ be a scaled $\mathrm{Cu}$-semigroup satisfying (O5) and (O7). Assume that $S$ has locally bounded comparison amplitude, and also that is upper and lower functionally almost divisible. Then $S$ is pure.
Proof.
We first prove that $S$ is almost unperforated. Let $x,y\in S$, $n\in{\mathbb{N}}$, and assume that $(n+1)x\leq ny$. We must show that $x\leq y$. This is equivalent to showing that $x^{\prime}\leq y$ for any $x^{\prime}\in S$ such that $x^{\prime}\ll x$.
Therefore, let $x^{\prime}\ll x$ and pick $x^{\prime\prime}\in S$ such that $x^{\prime}\ll x^{\prime\prime}\ll x$. Since we have $\widehat{x}\leq\tfrac{n}{n+1}\widehat{y}$, we may choose $\gamma,\gamma^{\prime}$ such that $\tfrac{n}{n+1}<\gamma<\gamma^{\prime}<1$. Using Proposition 2.3 (i) one gets $\widehat{x^{\prime\prime}}\ll\gamma\widehat{y}$, which allows us to pick $y^{\prime}\in S$ such that
$\widehat{x^{\prime\prime}}\ll\gamma\widehat{y^{\prime}},\,\,\,\text{ and }\,\,\,y^{\prime}\ll y.$
Take $y^{\prime\prime}$ such that $y^{\prime}\ll y^{\prime\prime}\ll y$, and choose $d\in{\mathbb{N}}$ such that $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$. Since $(S,\Sigma)$ has locally bounded comparison amplitude, we obtain $N=N(\gamma^{\prime},d)\in{\mathbb{N}}$ such that
$\widehat{v}\leq\gamma^{\prime}\widehat{w}\quad\Rightarrow\quad v\leq_{N}w,\quad\text{ for all }v,w\in\Sigma^{(d)}.$
Let $M\in{\mathbb{N}}$ be such that $M\geq 2$, $M\geq N$ and $\gamma\tfrac{M+1}{M-1}\leq\gamma^{\prime}$. Using (UFAD) for $x^{\prime}\ll x^{\prime\prime}$ and $M-1$, we obtain $v\in S$ satisfying
$v\leq x^{\prime\prime},\quad(M-1)\widehat{v}\leq\widehat{x^{\prime\prime}},\,\,\,\text{ and }\,\,\,x^{\prime}\leq Mv.$
Applying (LFAD) for $y^{\prime}\ll y^{\prime\prime}$ and $M$, one gets $w\in S$ such that
$Mw\leq y^{\prime\prime},\,\,\,\text{ and }\,\,\,\widehat{y^{\prime}}\leq(M+1)\widehat{w}.$
Combining the functional inequalities for $\widehat{v},\widehat{w}$ and using that $\widehat{x^{\prime\prime}}\leq\gamma\widehat{y^{\prime}}$ at the second step, we have
$(M-1)\widehat{v}\leq\widehat{x^{\prime\prime}}\leq\gamma\widehat{y^{\prime}}\leq\gamma(M+1)\widehat{w},$
which implies, by our choice of $M$, that $\widehat{v}\leq\gamma^{\prime}\widehat{w}$.
Note that both $v$ and $w$ belong to $\Sigma^{(d)}$, since $v\leq x^{\prime\prime},Mw\leq y^{\prime\prime}$, and $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$. We deduce that $v\leq_{N}w$ and, since $M\geq N$, we get $Mv\leq Mw$.
Hence,
$x^{\prime}\leq Mv\leq Mw\leq y^{\prime\prime}\leq y,$
as desired.
To show that $S$ is almost divisible, let $k\geq 2$ and take $x^{\prime},x\in S$ such that $x^{\prime}\ll x$. Since $S$ satisfies (LFAD), we may apply Lemma 4.4(ii) to obtain $y\in S$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+\frac{1}{k}\right)\widehat{y}.$
It follows that $x^{\prime}\leq\infty y$; see, for example, [APRT22, Lemma 6.6 (iii)]. We also have that $\lambda(x^{\prime})<\lambda((k+1)y)$ for every functional normalized at $y$. Since $S$ is almost unperforated, we obtain $x^{\prime}\leq(k+1)y$ by [ERS11, Proposition 6.2]. ∎
Theorem 5.7.
Let $A$ be a $C^{*}$-algebra that has locally bounded comparison amplitude and is $n$-almost divisible for some $n\in{\mathbb{N}}$. Then $A$ is pure.
Proof.
Since $\operatorname{Cu}(A)$ is $n$-almost divisible, Theorem 4.7 gives that $\operatorname{Cu}(A)$ has (LFAD) and (UFAD), and the result follows from Proposition 5.6. ∎
Theorem 5.8.
If a $C^{*}$-algebra is $(m,n)$-pure for some $m,n\in{\mathbb{N}}$, then it is pure.
Proof.
If $A$ is $(m,n)$-pure, then since it has $m$-comparison Lemma 3.6 implies that $A$ has locally bounded comparison amplitude. Since it is also $n$-almost divisible, we may apply Theorem 5.7 to conclude that $A$ is pure. ∎
6. C*-algebras with the Global Glimm Property and finite nuclear dimension
The next result is our variation of [RT17, Proposition 3.2(ii)] for nowhere scattered $C^{*}$-algebras. This has essentially appeared in the argument of [Vil23, Proposition 4.1].
Lemma 6.1.
Let $A$ be a nowhere scattered $C^{*}$-algebra with $\dim_{\mathrm{nuc}}(A)\leq m$. Then for every $k\geq 1$, and every pair $x^{\prime},x\in\operatorname{Cu}(A)$ satisfying $x^{\prime}\ll x$, there exists $y\in\operatorname{Cu}(A)$ such that
$y\ll x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll ky\ll 2(m+1)x.$
Proof.
By [TV24, Proposition 4.12] and [WZ10, Proposition 2.3], $A\otimes\mathcal{K}$ is nowhere scattered and of nuclear dimension at most $m$. We may thus assume that $A$ is stable.
Choose $a\in A_{+}$ and $\varepsilon>0$ such that
$x^{\prime}\ll[(a-\varepsilon)_{+}],\,\,\,\text{ and }\,\,\,[a]\ll x.$
Since $A$ is nowhere scattered, the hereditary sub-$C^{*}$-algebra $\overline{aAa}$ has no finite-dimensional representations, by [TV24, Theorem 3.1]. It also has nuclear dimension at most $m$ (by [WZ10, Proposition 2.5]), whence we can apply [RT17, Proposition 3.2(ii)] for $k$ (and for $\overline{aAa}$, $a$ and $\varepsilon$) to obtain $b\in\overline{aAa}_{+}$ such that
$[(a-\varepsilon)_{+}]\leq k[b]\leq 2(m+1)[a].$
Then $y:=[b]\in\operatorname{Cu}(A)$ has the desired properties. ∎
The following is inspired by [RT17, Lemma 3.4].
Lemma 6.2.
Let $A$ be a $C^{*}$-algebra with the Global Glimm Property, and let $m,l\in{\mathbb{N}}$. Assume that $\dim_{\mathrm{nuc}}A\leq m$. Then, there exists $L=L(m,l)$ such that, for every $x^{\prime}\ll x$, there exist $y_{0},\ldots,y_{l}\in\operatorname{Cu}(A)$ such that
$y_{0}+\ldots+y_{l}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll Ly_{j}$
for $j=0,\ldots,l$.
Proof.
Let $x^{\prime},x\in\operatorname{Cu}(A)$ be such that $x^{\prime}\ll x$. We claim that there exists a separable sub-$C^{*}$-algebra $B\subseteq A$ such that
  • (i)
    $B$ has the Global Glimm Property;
  • (ii)
    $\dim_{\mathrm{nuc}}(B)\leq m$;
  • (iii)
    the induced inclusion $\operatorname{Cu}(\iota)\colon\operatorname{Cu}(B)\to\operatorname{Cu}(A)$ is an order-embedding;
  • (iv)
    $x^{\prime}$ and $x$ are in the image of $\operatorname{Cu}(\iota)$.
Indeed, we know from Proposition 2.3(iii) and 2.6 in [WZ10] that the property ‘$\dim_{\mathrm{nuc}}(\_\,)\leq m$’ is separably inheritable. Using the same methods from [TV21, Section 5], one can also see that the Global Glimm Property satisfies the same condition. Now [TV21, Proposition 6.1] allows us to choose a sub-$C^{*}$-algebra $B$ with the required conditions.
Therefore we may assume that $A$ is separable and that, as in the proof of Lemma 6.1 above, we may also assume that $A$ is stable.
Now we apply induction on $l$. Note that, if $l=0$, one can take $y_{0}:=x$. Let us assume that $l=1$.
Choose $x^{\prime\prime}\in\operatorname{Cu}(A)$ such that $x^{\prime}\ll x^{\prime\prime}\ll x$. Since $A$ has the Global Glimm Property, $\operatorname{Cu}(A)$ is $(2,\omega)$-divisible by [TV23, Theorem 3.6]. Applied for $x^{\prime\prime}$ and $x$, we obtain $c\in\operatorname{Cu}(A)$ such that
$2c\leq x,\,\,\,\text{ and }\,\,\,x^{\prime\prime}\ll\infty c.$
Choose $c^{\prime}\in\operatorname{Cu}(A)$ such that $x^{\prime\prime}\leq\infty c^{\prime}$ and $c^{\prime}\ll c$. Using Lemma 2.1 with $c^{\prime}\ll c$ and $2c\leq x$, we find $d\in\operatorname{Cu}(A)$ such that
$c^{\prime}+d\leq x\leq 2d.$
Set $e:=d\wedge\infty c^{\prime}$, which is possible by Theorem 2.2. Using at the first step that $x^{\prime\prime}\leq\infty c^{\prime}$ and $x^{\prime\prime}\leq x\leq 2d$, we get
$x^{\prime\prime}\leq(2d)\wedge\infty c^{\prime}=2(d\wedge\infty c^{\prime})=2e.$
Choose $e^{\prime\prime},e^{\prime}\in\operatorname{Cu}(A)$ such that
$x^{\prime}\ll 2e^{\prime\prime},\,\,\,\text{ and }\,\,\,e^{\prime\prime}\ll e^{\prime}\ll e.$
Applying Lemma 6.1 for $k=2m+3$ and $e^{\prime\prime}\ll e^{\prime}$, we obtain $f\in\operatorname{Cu}(A)$ such that
(2) $f\ll e^{\prime},\,\,\,\text{ and }\,\,\,e^{\prime\prime}\ll(2m+3)f\ll 2(m+1)e^{\prime}.$
Choose $f^{\prime}\in\operatorname{Cu}(A)$ such that
$e^{\prime\prime}\ll(2m+3)f^{\prime},\,\,\,\text{ and }\,\,\,f^{\prime}\ll f.$
Applying (O5) for $f^{\prime}\ll f\leq e^{\prime}$, we obtain $g\in\operatorname{Cu}(A)$ such that
(3) $f^{\prime}+g\leq e^{\prime}\leq f+g.$
Set $y_{0}:=f^{\prime}$ and $y_{1}:=g+c^{\prime}$. Then
$y_{0}+y_{1}=f^{\prime}+g+c^{\prime}\leq e^{\prime}+c^{\prime}\leq d+c^{\prime}\leq x.$
Further,
$x^{\prime}\ll 2e^{\prime\prime}\ll 2(2m+3)f^{\prime}=(4m+6)y_{0}.$
Multiplying (3) by $32m+$ 3 at the first step, and using (2) at the second step, we deduce
(4) $(2m+3)e^{\prime}\leq(2m+3)f+(2m+3)g\leq(2m+1)e^{\prime}+(2m+3)g.$
We claim that $\widehat{e^{\prime}}\leq(2m+3)\widehat{y_{1}}$. To show this, let $\lambda\in F(S)$. We distinguish two cases:
First, assume that $\lambda(c^{\prime})<\infty$. Using that $e^{\prime}\ll e\leq\infty c^{\prime}$, it follows that $\lambda(e^{\prime})<\infty$ which allows, in (4), to cancel after applying $\lambda$ to get
$\lambda(e^{\prime})\leq 2\lambda(e^{\prime})\leq(2m+3)\lambda(g)\leq(2m+3)\lambda(y_{1}).$
On the other hand, if $\lambda(c^{\prime})=\infty$, then
$\lambda(e^{\prime})\leq\infty=(2m+3)\lambda(c^{\prime})\leq(2m+3)\lambda(y_{1}).$
Since $\operatorname{Cu}(A)$ has $m$-comparison (see Remarks 3.2 (1)), we get
$e^{\prime}\leq 2m(2m+3)y_{1}.$
Using that $x^{\prime}\ll 2e^{\prime}$, one has
$x^{\prime}\ll 2e^{\prime}\leq 4m(2m+3)y_{1}.$
Thus, setting $L:=\max\{4m+6,4m(2m+3)\}$, we see that $y_{0}+y_{1}\leq x$ and $x^{\prime}\ll Ly_{0},Ly_{1}$. This shows the statement for $l=1$.
Now let $l\in{\mathbb{N}}$ and assume that we have proven the statement for $l-1$.
Take $x^{\prime},x\in\operatorname{Cu}(A)$ such that $x^{\prime}\ll x$. We get $L^{\prime}\in{\mathbb{N}}$ and elements $z_{0},\ldots,z_{l-1}$ such that
$z_{0}+\ldots+z_{l-1}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll L^{\prime}z_{j}$
for $j=0,\ldots,l-1$.
Set $y_{j}:=z_{j}$ for each $j\leq l-2$, and take $z_{l-1}^{\prime}\in\operatorname{Cu}(A)$ such that $z_{l-1}^{\prime}\ll z_{l-1}$ and $x^{\prime}\ll L^{\prime}z_{l-1}^{\prime}$. By the argument above, one finds $L^{\prime\prime}\in{\mathbb{N}}$ and $y_{l-1},y_{l}\in\operatorname{Cu}(A)$ satisfying
$y_{l-1}+y_{l}\leq z_{l-1},\,\,\,\text{ and }\,\,\,z_{l-1}^{\prime}\ll L^{\prime\prime}y_{l-1},L^{\prime\prime}y_{l}.$
One can now check that the elements $y_{0},\ldots,y_{l}$ and $L:=L^{\prime}L^{\prime\prime}$ satisfy the required conditions. ∎
Proposition 6.3.
Let $A$ be a $C^{*}$-algebra with the Global Glimm Property, and let $m\in{\mathbb{N}}$. Assume that $\dim_{\mathrm{nuc}}A\leq m$. Then, there exists $N=N(m)$ such that $\operatorname{Cu}(A)$ is $N$-almost divisible.
Proof.
The proof is inspired by that of [RT17, Lemma 3.6]. Let $L$ be the constant obtained from Lemma 6.2 for $m$ and $l=m$. We verify the statement for $N:=3L(m+1)$.
Let $x^{\prime}\ll x$ and $k\geq 1$. We need to find $y$ such that
$ky\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)(N+1)y.$
Applying Lemma 6.1 for $3kL(m+1)$ and $x^{\prime}\ll x$, we obtain $c\in\operatorname{Cu}(A)$ such that
$c\ll x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll 3kL(m+1)c\ll 2(m+1)x.$
Choose $c^{\prime}\in\operatorname{Cu}(A)$ such that
$x^{\prime}\ll 3kL(m+1)c^{\prime},\,\,\,\text{ and }\,\,\,c^{\prime}\ll c.$
Now choose $x^{\prime\prime}\in\operatorname{Cu}(A)$ such that
$3kL(m+1)c^{\prime}\ll 2(m+1)x^{\prime\prime},\,\,\,\text{ and }\,\,\,x^{\prime\prime}\ll x.$
Applying Lemma 6.2 for $x^{\prime\prime}\ll x$, we obtain $y_{0},\ldots,y_{m}$ such that
$y_{0}+\ldots+y_{m}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime\prime}\leq Ly_{j}\quad\text{ for $j=0,\ldots,m$}.$
For each $j$, we obtain
$3kL(m+1)c^{\prime}\ll 2(m+1)x^{\prime\prime}\ll 2L(m+1)y_{j}.$
and thus $k\widehat{c^{\prime}}<_{s}\widehat{y_{j}}$.
By $m$-comparison, we get
$kc^{\prime}\leq y_{0}+\ldots+y_{m}\leq x.$
Further,
$x^{\prime}\ll 3kL(m+1)c^{\prime}\leq 3(k+1)L(m+1)c^{\prime}=(k+1)Nc^{\prime}\leq(k+1)(N+1)c^{\prime},$
which shows that $y=c^{\prime}$ has the desired properties. ∎
Theorem 6.4.
Every $C^{*}$-algebra with the Global Glimm Property and finite nuclear dimension is pure.
Proof.
Let $A$ be a $C^{*}$-algebra with the Global Glimm Property and that has nuclear dimension at most $m$. By Remarks 3.2 (1), we have that $A$ has $m$-comparison. Using Proposition 6.3, we know that $A$ is $n$-almost divisible for some $n$. Hence, $A$ is $(m,n)$-pure, and the result now follows from Theorem 5.8. ∎
Remark 6.5.
In [RT17, Theorem 3.1] it is shown that if a nowhere scattered $C^{*}$-algebra $A$ has finite nuclear dimension and no quotient of $A$ has a simple, purely infinite ideal, then $A$ is pure. Further, every approximately subhomogeneous $C^{*}$-algebra (ASH-algebra) has locally finite nuclear dimension [NW06]. By [ENST20, Theorem A], every $\mathcal{Z}$-stable ASH-algebra has decomposition rank at most two, and therefore finite nuclear dimension.
Combining these known results, we obtain another partial positive answer to Question B: For a separable, nowhere scattered ASH-algebra whose primitive ideal space has a basis of compact-open subsets or is Hausdorff, the properties ‘finite nuclear dimension’, ‘$\mathcal{Z}$-stability’ and ‘pureness’ are equivalent. If one additionally assumes that the algebra has stable rank one, then these properties are also equivalent to ‘strict comparison of positive elements’.

### v3

@@
 Abstract.
-We show that every $C^{*}$-algebra that is $(m,n)$-pure in the sense of Winter is already pure, that is, its Cuntz semigroup is almost unperforated and almost divisible.
-More generally, we show that even weaker comparison and divisibility properties automatically lead to pureness.
-As an application to the non-simple Toms-Winter conjecture, we show that every $C^{*}$-algebra with the Global Glimm Property and finite nuclear dimension is pure.
+We demonstrate that pure $C^{*}$-algebras form a robust class by proving that pureness follows from very weak comparison and divisibility properties.
+Using this, we show that every simple, non-elementary $C^{*}$-algebra with a unique quasitrace and with very mild comparison is pure, and, as a result, has strict comparison.
+Furthermore, sufficiently non-commutative $C^{*}$-algebras of stable rank one and with weak comparison are likewise pure.
+We also show that adequately non-elementary $C^{*}$-algebras with finite nuclear dimension are pure, which leads to the verification of the non-simple Toms-Winter conjecture for a large class of $C^{*}$-algebras.
 Key words and phrases:
@@
 Primary 46L05; Secondary 19K14, 46L80, 46L85.
-RA and FP were partially supported by Spanish Research State Agency (grant No.
-PID2020-113047GB-I00/AEI/10.13039/501100011033), by the Comissionat per Universitats i Recerca de la Generalitat de Catalunya (grant No.
-2021-SGR-01015) and by the Spanish State Research Agency through the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020-001084-M).
-HT was partially supported by the Knut and Alice Wallenberg Foundation (KAW 2021.0140).
+RA, FP and EV were partially supported by the Spanish Research State Agency (grant No.
+PID2020-113047GB-I00/AEI/10.13039/501100011033) and by the Comissionat per Universitats i Recerca de la Generalitat de Catalunya (grant No.
+2021-SGR-01015).
+RA and FP were also partially supported by the Spanish State Research Agency through the Severo Ochoa and María de Maeztu Program for Centers and Units of Excellence in R&D (CEX2020-001084-M).
+HT and EV were partially supported by the Knut and Alice Wallenberg Foundation (KAW 2021.0140).
+EV was also supported by the Fields Institute for Research in Mathematical Sciences.
 1. Introduction
-The groundbreaking classification of amenable factors by Connes and Haagerup [Con76, Haa87] in the 1970s and 80s sparked the Elliott program to classify simple, nuclear $C^{*}$-algebras.
-A first breakthrough in this direction was obtained with the classification in the purely infinite case by Kirchberg and Phillips [Kir95, Phi00] in the 1990s.
-In the stably finite case, the game-changing examples of Toms [Tom08] showed that additional regularity properties have to be assumed for $K$-theoretic classification, and building on decades of research by numerous mathematicians, in a series of spectacular results [Win14, EGLN15, TWW17, GLN20], the classification of simple, nuclear $C^{*}$-algebras that are $\mathcal{Z}$-stable and satisfy the universal coefficient theorem (UCT) was accomplished in the 2010s.
-(See also the recent approach [CGS+23] and the survey [Whi23].)
-While the UCT holds in all cases of interest (and possibly it is automatic for all nuclear $C^{*}$-algebras), it is known that $\mathcal{Z}$-stability can fail for simple, nuclear $C^{*}$-algebras, and it becomes highly relevant to determine when precisely this regularity property holds.
-To investigate this problem, Winter [Win12] introduced the notion of (m,n)-pureness for simple $C^{*}$-algebras, with $m$ and $n$ quantifying certain comparison and divisibility properties among operators, and he showed that every simple $C^{*}$-algebra that has locally finite nuclear dimension (a strong form of nuclearity) and that is $(m,n)$-pure for some $m,n$ is $\mathcal{Z}$-stable.
-Building on work of Rørdam [Rør04], Winter also showed that every simple, $\mathcal{Z}$-stable $C^{*}$-algebra is $(0,0)$-pure (abbreviated to just pure), which means that it satisfies the strongest of the comparison and divisibility properties.
-The concept of $(m,n)$-pureness does not depend on the ideal structure of the algebra, hence may be stated for non-simple $C^{*}$-algebras; see Definition 5.1.
-However, as we show in this paper, all versions of pureness coincide.
-Our first main result is (see Theorem 5.8):
-Theorem A.
-If a $C^{*}$-algebra is $(m,n)$-pure for some $m,n\in{\mathbb{N}}$, then it is pure.
-This should be contrasted with the well-known examples of simple, amenable $C^{*}$-algebras that do not have 00-comparison ([Vil98]), or that do not have 00-almost divisibility ([DHTW09]).
-Thus Theorem A can be interpreted as saying that if a $C^{*}$-algebra is not pure, then it necessarily has poor comparison or poor divisibility properties.
-Previously, Theorem A has been obtained for separable, simple, unital $C^{*}$-algebras with locally finite nuclear dimension [Win12], and recently for arbitrary simple $C^{*}$-algebras [APRT24, Theorem 10.5].
-Our proof of the general non-simple, non-nuclear case is based on the methods developed in [APRT24] in combination with new techniques about what we term functional divisibility (see Section 4 for details).
-In turn, these are partly inspired by results in [RT17].
-Thus a key ingredient in our approach is the Cuntz semigroup and its technical features (see Section 2 for the precise definition).
-This is an invariant constructed by Cuntz in [Cun78], which is modelled after the Murray-von Neumann semigroup of projections but using positive elements instead.
-The version we use is what is commonly called the complete Cuntz semigroup, considered in [CEI08], and technically more suitable as it behaves well with respect to, for example, inductive limits.
-The power of the Cuntz semigroup was demonstrated in the context of the classification program: A.
-S.
-Toms produced in [Tom08] two $C^{*}$-algebras that were indistinguishable by the Elliott invariant and a number of reasonable extensions of it, but could be told apart using their Cuntz semigroup.
-The salient distinguishing feature was strict comparison of positive elements.
-Another, more recent use of this semigroup is related to the structure of $C^{*}$-algebras of stable rank one, carried out in [APRT22], where powerful Cuntz semigroup techniques yielded solutions to three structural open problems.
-The Toms-Winter conjecture
-This prominent conjecture [Win18] aims at establishing a robust class of $C^{*}$-algebras covered by the classification theorem referred to above.
-It predicts that for a separable, simple, unital, non-elementary, nuclear $C^{*}$-algebra, the following properties are equivalent:
+Pureness is a regularity property for $C^{*}$-algebras introduced by Winter in his seminal investigation into $\mathcal{Z}$-stability and finite nuclear dimension of simple, nuclear $C^{*}$-algebras [Win12].
+Interest in these concepts was sparked by Toms’ groundbreaking examples [Tom08], where he constructed two simple, nuclear $C^{*}$-algebras with the same $K$-theoretic and tracial data —one pure, $\mathcal{Z}$-stable (i.e., tensorially absorbing the Jiang-Su algebra $\mathcal{Z}$), and of finite nuclear dimension, and the other not.
+This led to a revision of the Elliott program, aiming at classifying separable, simple, nuclear $C^{*}$-algebras.
+Building on decades of research by numerous mathematicians, this was ultimately achieved in a series of remarkable results [Win14, TWW17, GLN20].
+(See also the recent approach [CGS+23] and the survey [Whi23].) The classification theorem states that separable, simple, nuclear $C^{*}$-algebras that are $\mathcal{Z}$-stable and satisfy the universal coefficient theorem (UCT) can be distinguished up to isomorphism by their Elliott invariant, which essentially consists of topological K-Theory and the tracial simplex.
+As defined by Winter, a $C^{*}$-algebra is pure if it has strict comparison and is almost divisible.
+These are $C^{*}$-algebraic analogs of two fundamental properties of projections in $\mathrm{II}_{1}$ factors discovered by Murray and von Neumann: first, the comparison of projections in a $\mathrm{II}_{1}$ factor is determined by its unique trace $\tau$.
+Second, every projection $p$ is divisible in the sense that for every $n\geq 2$ there exists a projection $q$ with $\tau(q)=\tfrac{1}{n}\tau(p)$.
+Strict comparison for $C^{*}$-algebras was first considered by Blackadar [Bla88], and is currently a well-established concept in the structure theory of $C^{*}$-algebras.
+This property is difficult to prove and its validty has far-reaching implications (see, for example, its role in the Toms-Winter conjecture below 1.2).
+In fact, while comparison and divisibility hold automatically in every $\mathrm{II}_{1}$ factor, both strict comparison and almost divisibility can fail in simple $C^{*}$-algebras [Vil98, Vil99, DHTW09].
+With these parallels in mind, one can think of finite, simple, pure $C^{*}$-algebras as $C^{*}$-algebraic analogs of $\mathrm{II}_{1}$ factors.
+The concept also makes sense for infinite (that is, not finite) $C^{*}$-algebras, and it turns out that a simple $C^{*}$-algebra is purely infinite if and only if it is infinite and pure, and one can think of such algebras as the $C^{*}$-algebraic analogs of $\mathrm{III}$ factors.
+More generally, nonsimple pure $C^{*}$-algebras may be considered as analogs of von Neumann algebras with trivial type $\mathrm{I}$ summand.
+Nowadays, the importance of pureness goes beyond its original intent.
+For separable $C^{*}$-algebras, pureness of the central sequence algebra [KR14] is equivalent to $\mathcal{Z}$-stability of the underlying $C^{*}$-algebra ([PTV24]); for non-nuclear $C^{*}$-algebras —such as those coming from non-amenable groups or certain infinite reduced free products— pureness becomes an instrument weaker than $\mathcal{Z}$-stability for establishing strict comparison; for a $C^{*}$-algebra with mild divisibility (resp.
+comparison) properties, showing that it is not pure implies that is has very poor comparison (resp.
+divisibility) in the precise sense of Theorem C below.
+Due to these and other applications, the study of pureness has produced many important open questions: Is there a non-simple version of the Toms-Winter conjecture (1.3)?
+Do all $\mathrm{C}^{*}$-simple groups have strict comparison?
+(Equivalently, are they all pure?
+It is known that many are, as shown in [AGKP24], but in general the question is open.) Are there real rank zero versions of Rørdam’s example [Rør03]?
+(Conversely, are all non-elementary, simple, real rank zero $C^{*}$-algebras pure?)
+It is thus of utmost importance to understand which $C^{*}$-algebras are pure.
+In this paper, we provide a new approach to the study of pureness by analyzing two new (and considerably weak) conceptual properties which we term controlled comparison and functional divisibility.
+The study of these notions —discussed in 1.1 below— allows us, among other things, to prove robustness results for the class of pure $C^{*}$-algebras.
+One of the main applications of this new approach is the uncovering of new families where pureness can be verified.
+(The notions of nowhere scatteredness and the Global Glimm property mentioned below can be thought of as suitable generalizations of non-elementariness to the non-simple case; see the discussion before; see the discussion before Question D.)
+Theorem A (5.12, 5.13).
+Let $A$ be a $C^{*}$-algebra with controlled comparison. Moreover, assume that
+  1. (i)
+    $A$ is nowhere scattered with real rank zero or stable rank one, or
+  2. (ii)
+    $A$ is simple, unital, non-elementary, with a unique quasitracial state.
+Then $A$ is pure. In particular, $A$ has strict comparison.
+Theorem B (6.5).
+Every $C^{*}$-algebra with the Global Glimm Property and finite nuclear dimension is pure.
+Theorem B naturally fits in the context of the non-simple Toms-Winter conjecture (1.3), while Theorem A can be applied to large families of $C^{*}$-algebras: The conditions in (ii) are satisfied by any exact $\mathrm{C}^{*}$-simple group ([BKKO17]) and by Villadsen algebras of the second type [Vil99], while (i) is known to hold for all Villadsen algebras of the first type [Vil98].
+Thus, exact $\mathrm{C}^{*}$-simple groups and Villadsen algebras have strict comparison if and only if they have controlled comparison, if and only if they are pure.
+As previously stated, it is an important open problem to determine if all non-elementary simple real rank zero $C^{*}$-algebras have strict comparison (or if they are $\mathcal{Z}$-stable under the assumption of nuclearity and separability).
+Part (i) of Theorem A also applies to this class of algebras, reducing the problem to proving controlled comparison.
+1.1. Controlled comparison and functional divisibility
+Together with his notion of pureness, Winter introduced in [Win12] the quantified versions of $(m,n)$-pureness, where $m$ and $n$ are non-negative integers measuring how big the failure of comparison ($m$) and divisibility ($n$) is.
+Surprisingly, it follows from Winter’s work that for simple $C^{*}$-algebras with locally finite nuclear dimension there is a dimension reduction phenomenon: $(m,n)$-pureness automatically implies $(0,0)$-pureness, that is, pureness.
+This was recently extended to the simple (possibly non-nuclear) setting in [APRT24, Theorem 10.5].
+This dimension reduction phenomenon is one of the main techniques for establishing pureness or strict comparison outside the $\mathcal{Z}$-stable case, and a consequence of our Theorem C is that it holds for all $C^{*}$-algebras.
+The novelty of this paper lies in the study of two unquantified notions which we call controlled comparison and functional divisibility.
+Each of these notions is weaker than any of their quantified counterparts, thus making them easier to check in concrete examples.
+Inspired by the study initiated in [RT17], we show that there is a beautiful interplay between these functional notions and relative central sequences.
+This ultimately allows us to bypass any ideal-lattice obstruction and obtain our main tool for verifying pureness:
+Theorem C (5.7).
+Given a $C^{*}$-algebra $A$, the following statements are equivalent:
+  1. (i)
+    $A$ is pure, that is, it has strict comparison and is almost divisible.
+  2. (ii)
+    $A$ has controlled comparison and is functionally divisible.
+Theorem C implies, in particular, that the dimension reduction phenomenon alluded to above holds in full generality: If a $C^{*}$-algebra is $(m,n)$-pure for some $m$ and $n$, then it is pure.
+The present result provides a general statement that can be applied to non-simple, non-nuclear, $C^{*}$-algebras such as, for example, central sequence algebras.
+Even in the simple case, Theorem C represents a substantial improvement of [APRT24, Theorem 10.5]: Given a $C^{*}$-algebra $A$ satisfying condition (i) or (ii) in Theorem A, it is not known if $A$ satisfies the divisibility property in $(m,n)$-pureness (known as $n$-almost divisibility) for any $n$.
+However, we show in Propositions 4.10 and 4.11 that any $C^{*}$-algebra satisfying condition (i) or (ii) in Theorem A is automatically functionally divisibile.
+1.2. The Toms-Winter conjecture
+This conjecture [Win18] aims at establishing a class of $C^{*}$-algebras covered by the classification theorem mentioned at the beginning of the introduction.
+Explicitly, it predicts that for a separable, simple, unital, non-elementary, nuclear $C^{*}$-algebra, the following properties are equivalent:
   1. (1)
-    $A$ has finite nuclear dimension.
+    The $C^{*}$-algebra has finite nuclear dimension, a noncommutative generalization of finite covering dimension introduced in [WZ10].
   2. (2)
-    $A$ is $\mathcal{Z}$-stable.
+    The $C^{*}$-algebra is $\mathcal{Z}$-stable.
   3. (3)
-    $A$ has strict comparison of positive elements.
-The nuclear dimension was introduced in [WZ10] as a noncommutative generalization of covering dimension.
-A $C^{*}$-algebra $A$ is said to be $\mathcal{Z}$-stable if it tensorially absorbs the Jiang-Su algebra $\mathcal{Z}$, that is, if $A\cong\mathcal{Z}\otimes A$.
-A $C^{*}$-algebra $A$ has strict comparison of positive elements if it has 00-comparison; see Definitions 3.1 and 5.1 for details.
-In two groundbreaking papers [Win12, CET+21], it was shown that conditions (1) and (2) of the conjecture are equivalent.
-This was later generalized to stably projectionless, simple $C^{*}$-algebras in [Tik14, CE20], thereby showing that a separable, simple, non-elementary $C^{*}$-algebra has finite nuclear dimension if and only if it is nuclear and $\mathcal{Z}$-stable.
+    The $C^{*}$-algebra has strict comparison of positive elements (see Section 3).
+In two groundbreaking papers [Win12, CET+21], it was shown that conditions (1) and (2) are equivalent.
 Further, it is known that (2) implies (3) [Rør04].
-The implication (3$\Rightarrow$(2) has been verified under certain additional assumptions [Sat12, KR14, TWW15, Thi20], but it remains open in general.
-The Toms-Winter conjecture is also closely related to pureness, which we view as a property lying between (2) and (3).
-Indeed, as mentioned above, $\mathcal{Z}$-stability implies pureness, which in turn implies strict comparison of positive elements; see Proposition 5.2.
-Further, among separable, simple, unital, non-elementary $C^{*}$-algebras, pureness implies $\mathcal{Z}$-stability under the additional assumption of locally finite nuclear dimension [Win12], and strict comparison of positive elements implies pureness under the additional assumption of a positive solution to the Rank Problem.
-Such a positive solution has been obtained for $C^{*}$-algebras of stable rank one [Thi20, APRT22].
-The non-simple Toms-Winter conjecture
-The regularity properties of the Toms-Winter conjecture are also relevant for non-simple $C^{*}$-algebras, and an investigation of their interplay was initiated in [RT17].
-While the assumption of unitality in the original formulation of the conjecture could be dispensed with, the condition to be non-elementary is crucial.
+The implication (3$\Rightarrow$(2) has been verified under certain additional assumptions [Sat12, KR14, TWW15, Thi20], but remains open in general.
+The Toms-Winter conjecture is closely related to pureness, which we view as a property lying between (2) and (3).
+Indeed, $\mathcal{Z}$-stability implies pureness, which in turn implies strict comparison of positive elements; see Proposition 5.2.
+Further, strict comparison of positive elements implies pureness under the additional assumption of a positive solution to the Rank Problem; see Section 2.
+Such a positive solution has been obtained for $C^{*}$-algebras of real rank zero [ER06] and of stable rank one [Thi20, APRT22].
+1.3. The non-simple Toms-Winter conjecture
+The regularity properties of the Toms-Winter conjecture are also relevant for non-simple $C^{*}$-algebras, and an investigation of their interplay was initiated by Robert and Tikuisis [RT17].
+While the assumption of unitality in the original formulation of the Toms-Winter conjecture could be dispensed with, the condition of being non-elementary is crucial.
 Indeed, elementary $C^{*}$-algebras have nuclear dimension zero, but they are neither $\mathcal{Z}$-stable nor pure.
-There are two natural generalizations of non-elementariness to the non-simple setting: Nowhere scatteredness, and the Global Glimm property.
-Here, a $C^{*}$-algebra is nowhere scattered if none of its quotients has a nonzero, elementary ideal, or equivalently if every hereditary sub-$C^{*}$-algebra is generated by nilpotent elements [TV24].
-A $C^{*}$-algebra has the Global Glimm Property if every hereditary sub-$C^{*}$-algebra contains an approximately full, nilpotent element [TV23].
-Clearly, every $C^{*}$-algebra with the Global Glimm Property is nowhere scattered, and the Global Glimm Problem asks whether the converse holds.
-We ask:
-Question B.
+The natural generalization of non-elementariness to the non-simple setting is the notion of nowhere scateredness: a $C^{*}$-algebra is nowhere scattered if none of its quotients has a nonzero, elementary ideal, or equivalently if every hereditary sub-$C^{*}$-algebra is generated by nilpotent elements [TV24a].
+This notion is tightly related to the Global Glimm Property: a $C^{*}$-algebra has the Global Glimm Property if every hereditary sub-$C^{*}$-algebra contains an approximately full, nilpotent element [TV23].
+Clearly, every $C^{*}$-algebra with the Global Glimm Property is nowhere scattered.
+The converse remains open, and is known as the Global Glimm Problem.
+In light of these considerations, the following question is natural:
+Question D.
 Let $A$ be a separable, nowhere scattered, nuclear $C^{*}$-algebra. Are the following properties equivalent?
@@
     $A$ has strict comparison of positive elements.
-The implications (2$\Rightarrow$(3a)$\Rightarrow$(3b) hold in general; see Proposition 5.2.
-We also note that pureness implies the Global Glimm Property; see [TV23, Theorem 3.6].
-Therefore, Question B contains in particular the question of whether the Global Glimm Problem has a positive solution for $C^{*}$-algebras with finite nuclear dimension.
-As for simple $C^{*}$-algebras, the question whether strict comparison of positive elements implies pureness is closely related to the Rank Problem.
-This was solved for $C^{*}$-algebras of real rank zero or stable rank one ([ER06] and [APRT22]), which shows that (3b) implies (3a) in these cases.
-In [RT17, Theorems 7.10, 7.15] it is shown that (3a) implies (2) under the additional assumption that the $C^{*}$-algebra $A$ has locally finite nuclear dimension, that no quotient of $A$ has a simple, purely infinite ideal, and that the primitive ideal space of $A$ has a basis of compact-open subsets or is Hausdorff.
-The following result (which is Theorem 6.4) provides a further partial answer to Question B.
-Theorem C.
-Every $C^{*}$-algebra with the Global Glimm Property and finite nuclear dimension is pure.
+The implications (2$\Rightarrow$(3a)$\Rightarrow$(3b) hold in general, and pureness always implies the Global Glimm Property; see Proposition 5.2.
+Therefore, Question D contains in particular the Global Glimm Problem for $C^{*}$-algebras with finite nuclear dimension.
+Partial results on the implication (3a)$\Rightarrow$(2) were established in [RT17, Theorems 7.10, 7.15], while the implication (2$\Rightarrow$(1) has been shown to hold in a variety of cases; see, for example, [BL24, BGSW22, ENST20].
+As for simple $C^{*}$-algebras, whether (3b) implies (3a) is related to the Rank Problem.
+Combining Theorem B with some of the above mentioned results, we obtain an answer to Question D for a relevant class of $C^{*}$-algebras, thereby verifying the non-simple Toms-Winter conjecture in this setting.
+Theorem E (6.6).
+Let $A$ be a separable, locally subhomogeneous $C^{*}$-algebra that has stable rank one, topological dimension zero, and the Global Glimm Property. Then, the answer to Question D is positive.
+Conventions
+Given a $C^{*}$-algebra, we let $A_{+}$ to denote its positive elements.
+We let ${\mathbb{N}}$ denote the set of natural numbers, including 00.
+Further, ${\mathcal{K}}$ denotes the $C^{*}$-algebra of compact operators on a separable, infinite-dimensional Hilbert space.
 Acknowledgements
-The authors thank Leonel Robert for valuable feedback on an earlier version of this paper.
+The authors thank Leonel Robert, Stuart White and Wilhelm Winter for valuable feedback on earlier versions of this paper.
 2. Preliminaries
-The techniques used in the paper mainly concern the structure of the Cuntz semigroup $\operatorname{Cu}(A)$ of a $C^{*}$-algebra $A$, whose definition and properties we describe below.
+The techniques used in this paper mainly concern the structure of the Cuntz semigroup of a $C^{*}$-algebra whose definition and properties we describe below.
 Given positive elements $a,b$ in a $C^{*}$-algebra $A$, one says that $a$ is Cuntz subequivalent to $b$, denoted by $a\precsim b$, if there is a sequence $(r_{n})_{n}$ in $A$ such that $\lim_{n\to\infty}\|a-r_{n}br_{n}^{*}\|=0$.
@@
 […]
-It is proved in [CEI08] that $\operatorname{Cu}(A)$ satisfies the following properties defined for a general positively ordered monoid $S$ (and usually referred to as axioms):
+It is proved in [CEI08] that $\operatorname{Cu}(A)$ satisfies the following properties (usually referred to as axioms):
   1. (O1)
-    If $(x_{n})_{n}$ is an increasing sequence in $S$, then $\sup_{n}x_{n}$ exists.
+    If $(x_{n})_{n}$ is an increasing sequence in $\operatorname{Cu}(A)$, then $\sup_{n}x_{n}$ exists.
   2. (O2)
-    For any $x\in S$ there exists a sequence $(x_{n})_{n}$ such that $x_{n}\ll x_{n+1}$ for all $n$ and $x=\sup_{n}x_{n}$. (We say that $(x_{n})_{n}$ is a $\ll$-increasing sequence.)
+    For any $x\in\operatorname{Cu}(A)$ there exists a sequence $(x_{n})_{n}$ such that $x_{n}\ll x_{n+1}$ for all $n$ and $x=\sup_{n}x_{n}$. (We say that $(x_{n})_{n}$ is a $\ll$-increasing sequence.)
   3. (O3)
-    If $x_{1}\ll x_{2}$ and $y_{1}\ll y_{2}$, then $x_{1}+y_{1}\ll x_{2}+y_{2}$.
+    If $x_{1}\ll x_{2}$ and $y_{1}\ll y_{2}$ in $\operatorname{Cu}(A)$, then $x_{1}+y_{1}\ll x_{2}+y_{2}$.
   4. (O4)
-    If $(x_{n})_{n}$ and $(y_{n})_{n}$ are increasing sequences in $S$, then $\sup_{n}(x_{n}+y_{n})=\sup_{n}x_{n}+\sup_{n}y_{n}$.
+    If $(x_{n})_{n}$ and $(y_{n})_{n}$ are increasing sequences in $\operatorname{Cu}(A)$, then $\sup_{n}(x_{n}+y_{n})=\sup_{n}x_{n}+\sup_{n}y_{n}$.
 The relation $\ll$ in these axioms is defined as follows: $x\ll y$ if for every increasing sequence $(y_{n})_{n}$ satisfying $y\leq\sup_{n}y_{n}$ there exists $n_{0}\in{\mathbb{N}}$ such that $x\leq y_{n_{0}}$.
-The relation $\ll$ is called the way-below relation, or compact containment relation, and one says that ‘$x$ is way-below $y$’ if $x\ll y$.
+The relation $\ll$ is called the way-below relation, or the compact containment relation, and one says that ‘$x$ is way-below $y$’ (or that ‘$x$ is compactly contained in $y$’) if $x\ll y$.
 An element $u\in S$ such that $u\ll u$ is termed compact.
-A positively ordered monoid satisfying axioms (O1)-(O4) is called a $\mathrm{Cu}$-semigroup and belongs to a category of semigroups that has thoroughly been studied (see, for example, [APT18], [APT20a], [APT20b] [APT20c], [TV22], [APRT22], among others; see also the recent survey [GP23]).
-Yet, when dealing with semigroups obtained as Cuntz semigroups of $C^{*}$-algebras, or further restricting to particular families of $C^{*}$-algebras, additional axioms and properties are satisfied.
-The first axiom that is added after (O4) is the following:
+A positively ordered monoid satisfying axioms (O1)-(O4) is called a $\mathrm{Cu}$-semigroup and belongs to a category of semigroups that has thoroughly been studied (see, for example, [APT18, APT20c, APT20a, APT20b], among others; see also the recent survey [GP24]).
+Yet, when dealing with semigroups obtained as Cuntz semigroups of $C^{*}$-algebras, additional axioms and properties are satisfied:
   1. (O5)
@@
 This axiom is usually referred to as the axiom of almost algebraic order.
-The first part of the statement above was proved in [RW10, Lemma 7.1], and the second part can be found in [APT18, Proposition 4.6].
+For the Cuntz semigroup of an arbitrary $C^{*}$-algebra, the first part of the statement above was proved in [RW10, Lemma 7.1], and the second part can be found in [APT18, Proposition 4.6].
 Since we will often use the following consequence of (O5), we add a proof here for completeness.
@@
 Recall that a $\mathrm{Cu}$-semigroup $S$ is said to be countably based if all elements can be written as suprema of elements from a fixed countable subset.
-It is known that $\operatorname{Cu}(A)$ is countably based whenever $A$ is separable (for a proof, see [APS11, Lemma 1.3]).
-Theorem 2.2 (cf. [APRT21, Theorems 2.4, 2.5]).
+It is known that $\operatorname{Cu}(A)$ is countably based whenever $A$ is separable; see [APS11, Lemma 1.3].
+Theorem 2.2 ([APRT21, Theorems 2.4, 2.5]).
 Let $S$ be a countably based $\mathrm{Cu}$-semigroup satisfying (O5)-(O7).
@@
 Given $x\in S$, we obtain a map $\widehat{x}\colon F(S)\to[0,\infty]$ defined by $\widehat{x}(\lambda)=\lambda(x)$, for $\lambda\in F(S)$.
-$\widehat{x}$ is linear and lower semicontiuous, and if $S$ satisfies (O5) and (O6), then $\textrm{Lsc}(F(S))$, the set of all such maps, is an algebraically ordered, lattice ordered, compact cone.
+The map $\widehat{x}$ is linear and lower semicontinuous.
+Denote by $\operatorname{Lsc}(F(S))$ the set of all linear, lower semicontinuous maps on $F(S)$ with values in $[0,\infty]$.
 Two main questions arise in this setting.
-To determine the range of the so-called rank map $x\mapsto\widehat{x}$ (referred to as the Rank Problem); and determine to what extent the order in $S$ can be inferred by the order in $\textrm{Lsc}(F(S))$.
-Note that $x\mapsto\widehat{x}$ is additive and preserves sequential suprema; it does not preserve the relation of compact containment, and is certainly not an order embedding.
-However, the result below will be useful.
+The first concerns the extent to which the ordering in $A$ can be inferred from the ordering in $\textrm{Lsc}(F(S))$.
+Note that whilst $x\mapsto\widehat{x}$ is an additive map that preserves the order and suprema of increasing sequences, it typically does not preserve the way-below relation and, moreover, is generally not an order embedding.
+Still, the result below will be useful.
 Given elements $x$ and $y$ in a partially ordered semigroup, one writes $x<_{s}y$ if there exists $k\in{\mathbb{N}}$ such that $(k+1)x\leq ky$.
-Item (ii) in the following lemma is a consequence of [Rob13, Proposition 2.2.2] or [APT18, Theorem 5.2.19].
 Proposition 2.3 ([Rob13, Lemma 2.2.5], [APRT24, Lemma 5.3]).
@@
   2. (ii)
-    If $\widehat{x}\leq\gamma\widehat{y}$ for some $\gamma\in(0,1)$, then $x^{\prime}<_{s}y$ for all $x^{\prime}\ll x$.
+    If $\widehat{x}\leq\gamma\widehat{y}$ for some $\gamma\in(0,1)$, then $x^{\prime}<_{s}y$ for all $x^{\prime}\in S$ with $x^{\prime}\ll x$.
+The second question that arises consists of determining the range of the map $x\mapsto\widehat{x}$.
+This is referred to as the Rank Problem and was explored in detail in [Thi20] and [APRT22].
+For a $\mathrm{Cu}$-semigroup $S$, let us denote by $\textrm{L}(F(S))$ the smallest subsemigroup of $\operatorname{Lsc}(F(S))$ closed under suprema of increasing sequences and containing all elements of the form $\tfrac{1}{n}\widehat{x}$ (see [Rob13, Section 3]).
+Recall that an element $x\in S$ is termed soft if for every $x^{\prime}\in S$ with $x^{\prime}\ll x$ there exists $t\in S$ such that $x^{\prime}+t\ll x$ and $x^{\prime}\ll\infty t$ (see [TV24b, Definition 4.3, Remark 4.4]; for $C^{*}$-algebras of stable rank one, this agrees with [APT18, Definition 5.3.1]).
+The following theorem summarizes the realization results from [APRT22], which will be required in Section 4.
+The existence of the map $\alpha$ is shown in [APRT22, Theorem 7.2], and the properties of $\alpha$ are proved in [APRT22, Proposition 7.4, Theorem 7.13, Proposition 8.2, Theorem 8.4, Corollary 8.5].
+Theorem 2.4 ([APRT22]).
+Let $A$ be a separable, nowhere scattered $C^{*}$-algebra with stable rank one.
+Then there is a natural, additive, order-preserving map $\alpha\colon\textrm{L}(F(\operatorname{Cu}(A)))\to\operatorname{Cu}(A)$ that preserves suprema of increasing sequences and such that $f=\widehat{\alpha(f)}$ for every $f\in\textrm{L}(F(\operatorname{Cu}(A)))$.
+The image of $\alpha$ consists of soft elements, and we have
+$\alpha(f+\widehat{x})=\alpha(f)+x$
+whenever $x\in\operatorname{Cu}(A)$ and $f\in\textrm{L}(F(\operatorname{Cu}(A)))$ satisfy $\widehat{x}\leq\infty f$.
 3. Variants of comparison
-We first recall the definition of $m$-comparison as considered by Toms and Winter in [TW09, Lemma 6.1], and later formalized by Ortega, Perera and Rørdam in [OPR12, Definition 2.8].
+In this section we deal with various comparison properties.
+In particular, we recall the definition of $m$-comparison (first considered in [TW09, Lemma 6.1] and later formalized in [OPR12, Definition 2.8]) and clarify the equivalence between 00-comparison (also known as almost unperforation) with the notion of strict comparison for $C^{*}$-algebras.
+We also recall the concept introduced in [APRT24, Definition 6.1], which in this paper is renamed as controlled comparison.
 Definition 3.1.
-Let $m\in{\mathbb{N}}$.
-A $\mathrm{Cu}$-semigroup $S$ has $m$-comparison if for all $x,y_{0},\ldots,y_{m}\in S$ satisfying $x<_{s}y_{j}$ for $j=0,\ldots,m$, we have $x\leq y_{0}+\ldots+y_{m}$.
-The lowest possible value for $m$-comparison is attained when $m=0$, which is commonly referred to as saying that $S$ is almost unperforated.
-We say that a $C^{*}$-algebra has $m$-comparison if its Cuntz semigroup does.
+Let $m\in{\mathbb{N}}$. A $\mathrm{Cu}$-semigroup $S$ has $m$-comparison if for all $x,y_{0},\ldots,y_{m}\in S$ satisfying $x<_{s}y_{j}$ for $j=0,\ldots,m$, we have $x\leq y_{0}+\ldots+y_{m}$.
+We say that a $C^{*}$-algebra has $m$-comparison if its Cuntz semigroup does. A $\mathrm{Cu}$-semigroup or $C^{*}$-algebra is said to be almost unperforated if it has 00-comparison.
 Remarks 3.2.
-(1) L. Robert showed in [Rob11, Theorem 1.3] that, if $A$ is a $C^{*}$-algebra with nuclear dimension at most $m$, then $A$ has $m$-comparison.
+(1) Robert showed in [Rob11, Theorem 1.3] that a $C^{*}$-algebra has $m$-comparison whenever it has nuclear dimension at most $m$.
 (2) Let $m\in{\mathbb{N}}$ and let $S$ be a $\mathrm{Cu}$-semigroup that has $m$-comparison. Then, $x<_{s}y$ implies $x\leq(m+1)y$, for all $x,y\in S$.
-Given $N\in{\mathbb{N}}$ and elements $x$ and $y$ in a partially ordered semigroup, we write $x\leq_{N}y$ if $nx\leq ny$ for all $n\in{\mathbb{N}}$ with $n\geq N$; see [APRT24, Notation 5.2] and also [APT18, Definition 5.6.1].
-To recall the definition of locally bounded comparison amplitude, we need the notion of a scale in a $\mathrm{Cu}$-semigroup.
-3.3Scales.
-Let $S$ be a $\mathrm{Cu}$-semigroup. A scale in $S$ is a subset $\Sigma\subseteq S$ such that
+The notion of strict comparison was first considered in [Bla88]. For $\mathrm{Cu}$-semigroups, the concept was considered in [ERS11, Section 6].
+Definition 3.3.
+A $\mathrm{Cu}$-semigroup $S$ has strict comparison if for all $x,y\in S$ satisfying $x\leq\infty y$ and $\lambda(x)<\lambda(y)$ for any $\lambda\in F(S)$ with $\lambda(y)=1$, one has $x\leq y$.
+A $C^{*}$-algebra $A$ has strict comparison of positive elements by quasitraces (specifically, $[0,\infty]$-valued, lower semicontinuous $2$-quasitraces) if its Cuntz semigroup does.
+3.4Strict comparison and almost unperforation.
+We take the opportunity here to clarify the equivalence between the concepts of strict comparison and almost unperforation.
+Let $S$ be a $\mathrm{Cu}$-semigroup, and let $x,y\in S$.
+The following conditions relating the comparison of $x,y$ and $\widehat{x},\widehat{y}$ were considered in [APT18]:
   1. (i)
-    is downward hereditary: if $x\leq y$ for some $x\in S$ and $y\in\Sigma$, then $x\in\Sigma$,
+    We have $x<_{s}y$, that is, there exists $k\in{\mathbb{N}}$ such that $(k+1)x\leq ky$.
   2. (ii)
-    is closed under suprema of increasing sequences, and
+    We have $\widehat{x}<_{s}\widehat{y}$, that is, there exists $\varepsilon\in(0,1)$ such that $\widehat{x}\leq(1-\varepsilon)\widehat{y}$.
+  3. (iii)
+    We have $x\leq\infty y$ and $\lambda(x)<\lambda(y)$ for any $\lambda\in F(S)$ with $\lambda(y)=1$.
+  4. (iv)
+    We have $x^{\prime}<_{s}y$ for every $x^{\prime}\in S$ with $x^{\prime}\ll x$.
+It was shown in [APT18, Theorem 5.2.18] that the implications (i)$\Rightarrow$(ii)$\Rightarrow$(iii)$\Rightarrow$(iv) always hold.
+Using this, it follows that the following conditions are equivalent for $S$.
+(This was first obtained in [ERS11, Proposition 6.2]; see also [APT18, Proposition 5.2.20].)
+  1. (i’)
+    $S$ is almost unperforated.
+  2. (ii’)
+    If $x,y\in S$ satisfy $\widehat{x}\leq(1-\varepsilon)\widehat{y}$ for some $\varepsilon\in(0,1)$, then $x\leq y$.
+  3. (iii’)
+    $S$ has strict comparison.
+Now, let $A$ be a $C^{*}$-algebra.
+There is a natural correspondence between functionals on $\operatorname{Cu}(A)$ and quasitraces on $A$.
+This is defined by sending a quasitrace $\tau\colon(A\otimes\mathcal{K})_{+}\to[0,\infty]$ to the functional $d_{\tau}\colon\operatorname{Cu}(A)\to[0,\infty]$ given by
+$d_{\tau}([a])=\lim_{n\to\infty}\tau(a^{1/n})$
+for $a\in(A\otimes\mathcal{K})_{+}$; see, for example, [ERS11, Proposition 4.2].
+Using this, it follows that $A$ has strict comparison if and only if for all $a,b\in(A\otimes\mathcal{K})_{+}$ such that $a$ belongs to the closed ideal generated by $b$, and such that $d_{\tau}([a])<d_{\tau}([b])$ for every quasitrace $\tau$ with $d_{\tau}([b])=1$, we have $a\precsim b$.
+It also follows from the above considerations that a $C^{*}$-algebra $A$ has strict comparison (of positive elements by quasitraces) if and only if it is almost unperforated.
+Some of the confusion surrounding the notion of strict comparison arises since it is applied in a more restrictive setting, for example by considering comparison of projections (instead of positive elements), or by considering comparison by traces (instead of quasitraces); see [NR16, Section 3] for some clarifications.
+Another source of confusion is that strict comparison is often used in the context of simple $C^{*}$-algebras, in which case every element is automatically contained in the closed ideal generated by a nonzero element.
+Therefore, a simple, unital $C^{*}$-algebra $A$ has strict comparison if and only if for all nonzero elements $a,b\in(A\otimes\mathcal{K})_{+}$ such that $d_{\tau}([a])<d_{\tau}([b])$ for every normalized quasitrace $\tau$, we have $a\precsim b$.
+For the definition below, we need to recall the notion of a scale in a $\mathrm{Cu}$-semigroup.
+3.5Scales.
+Let $S$ be a $\mathrm{Cu}$-semigroup. Following [APT20c, Definition 4.1], a scale in $S$ is a subset $\Sigma\subseteq S$ satisfying the following conditions:
+  1. (i)
+    $\Sigma$ is downward hereditary: If $x\leq y$ for some $x\in S$ and $y\in\Sigma$, then $x\in\Sigma$.
+  2. (ii)
+    $\Sigma$ is closed under suprema of increasing sequences.
   3. (iii)
     $\Sigma$ generates $S$ as an ideal, that is, for $x^{\prime},x\in S$ with $x^{\prime}\ll x$, there exist $y_{1},\ldots,y_{n}\in\Sigma$ such that $x^{\prime}\leq y_{1}+\ldots+y_{n}$.
-(See [APT20c, Definition 4.1]). We say that the pair $(S,\Sigma)$ is a scaled $\mathrm{Cu}$-semigroup.
-Given a scale $\Sigma$ and $d\in{\mathbb{N}}$, the $d$-fold amplification of $\Sigma$ is defined as $\Sigma^{(0)}:=\{0\}$, and
+We say that the pair $(S,\Sigma)$ is a scaled $\mathrm{Cu}$-semigroup.
+Given a scale $\Sigma$ and $d\in{\mathbb{N}}$, the $d$-fold amplification of $\Sigma$ is defined as $\Sigma^{(0)}:=\{0\}$, and for $d\geq 1$ as
 […]
-for $d\geq 1$.
 If $A$ is a $C^{*}$-algebra, then
@@
 $\Sigma_{A}^{(d)}=\big{\{}x\in\operatorname{Cu}(A):x\leq[a]\text{ for some $a\in M_{d}(A)_{+}$}\big{\}}.$
-We recall below the definition of locally bounded comparison amplitude, from [APRT24]. Useful reformulations of this condition will be given in Lemma 3.5.
-Definition 3.4 ([APRT24, Definition 6.1]).
-A scaled $\mathrm{Cu}$-semigroup $(S,\Sigma)$ has locally bounded comparison amplitude, in short (LBCA), if for every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $N=N(\gamma,d)$ such that the following holds:
+Given $N\in{\mathbb{N}}$ and elements $x$ and $y$ in a partially ordered semigroup, we write $x\leq_{N}y$ if $nx\leq ny$ for all $n\in{\mathbb{N}}$ with $n\geq N$.
+The following was termed ‘locally bounded comparison amplitude’ in [APRT24, Definition 6.1].
+Definition 3.6.
+A scaled $\mathrm{Cu}$-semigroup $(S,\Sigma)$ has controlled comparison if for every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $N=N(\gamma,d)$ such that the following holds:
 $\widehat{x}\leq\gamma\widehat{y}\quad\textrm{ implies }\quad x\leq_{N}y,\quad\text{ for all }x,y\in\Sigma^{(d)}.$
-A $C^{*}$-algebra will be said to have locally bounded comparison amplitude if $(\operatorname{Cu}(A),\Sigma_{A})$ has (LBCA).
-Lemma 3.5.
-Let $(S,\Sigma)$ be a scaled $\mathrm{Cu}$-semigroup. The following are equivalent:
-  1. (i)
-    $(S,\Sigma)$ has locally bounded comparison amplitude.
-  2. (ii)
-    For every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $N=N(\gamma,d)$ such that
-    $\widehat{x}\leq\gamma\widehat{y^{\prime}},\quad y^{\prime}\ll y\quad\text{ implies }\quad Nx\leq Ny,\qquad\text{for all }x,y^{\prime},y\in\Sigma^{(d)}.$
-  3. (iii)
-    For every $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$ there exists $M=M(\gamma,d)$ such that:
-    $\widehat{x}\leq\gamma\widehat{y}\quad\text{ implies }\quad(M+1)x\leq My,\qquad\text{for all }x,y\in\Sigma^{(d)}.$
-Proof.
-It is clear that (i) implies (ii).
-Let us show that (ii) implies (iii). Fix $\gamma\in(0,1)$ and $d\in{\mathbb{N}}$. Choose $\gamma^{\prime}$ and $n\in{\mathbb{N}}$ such that
-$\gamma<\gamma^{\prime},\quad\frac{n+1}{n}\gamma^{\prime}<1.$
-Let $N=N(\tfrac{n+1}{n}\gamma^{\prime},d(n+1))$ as provided by (ii). We claim that $M=Nn$ has the desired properties to verify (iii) (for $\gamma$ and $d$).
-To this end, let $x,y\in\Sigma^{(d)}$ satisfy $\widehat{x}\leq\gamma\widehat{y}$.
-Let $x^{\prime}\in S$ be such that $x^{\prime}\ll x$.
-By Proposition 2.3 (i) applied to $x^{\prime}\ll x$ and $1<\frac{\gamma^{\prime}}{\gamma}$, we have that $\widehat{x^{\prime}}\ll\gamma^{\prime}\widehat{y}$.
-Hence, there exists $y^{\prime}\ll y$ such that $\widehat{x^{\prime}}\leq\gamma^{\prime}\widehat{y^{\prime}}$.
-Then
-$(n+1)\widehat{x^{\prime}}\leq(n+1)\gamma^{\prime}\widehat{y^{\prime}}=\left(\frac{n+1}{n}\gamma^{\prime}\right)n\widehat{y^{\prime}}.$
-Since $(n+1)x^{\prime},ny^{\prime}$ and $ny$ belong to $\Sigma^{((n+1)d)}$, it follows from the choice of $N$ that $N(n+1)x^{\prime}\leq Nny$.
-Passing to the supremum over all $x^{\prime}\ll x$ we get that $N(n+1)x\leq Nny$, and thus $(M+1)x\leq My$.
-Finally, note that if elements $x$ and $y$ in a partially ordered semigroup satisfy $(M+1)x\leq My$ for some $M\in{\mathbb{N}}$, then there exists $N\in{\mathbb{N}}$ such that $(n+1)x\leq ny$ for all $n\geq N$, and in particular $x\leq_{N}y$; see [APT18, Proposition 5.2.13].
-This shows that (iii) implies (i).
-∎
-The following lemma is implied by [APRT24, Proposition 10.3].
-Lemma 3.6.
-Any $C^{*}$-algebra with $m$-comparison has locally bounded comparison amplitude.
+We say that a $C^{*}$-algebra $A$ has controlled comparison if $(\operatorname{Cu}(A),\Sigma_{A})$ does.
+The next result follows easily from [APRT24, Proposition 10.3].
+Proposition 3.7.
+If a $C^{*}$-algebra has $m$-comparison for some $m\in{\mathbb{N}}$, then it has controlled comparison.
 Proof.
 In order to apply [APRT24, Proposition 10.3] it is required that the Cuntz semigroup of a $C^{*}$-algebra satisfies (O5), (O6), and the so-called Edward’s condition.
-The first two are always satisfied, as mentioned in the preliminaries.
-As for the third condition, this was also shown to hold in general in [APRT21, Theorem 5.3].
+The first two are always satisfied, as mentioned in the preliminaries, and the third condition was shown to hold in [APRT21, Theorem 5.3].
 ∎
-Theorem 3.7.
-If a $C^{*}$-algebra has finite nuclear dimension, then its Cuntz semigroup has locally bounded comparison amplitude.
+Corollary 3.8.
+Every $C^{*}$-algebra with finite nuclear dimension has controlled comparison.
 Proof.
-Let $A$ be a $C^{*}$-algebra with nuclear dimension at most $m$ for some $m\in{\mathbb{N}}$.
-As observed in Remarks 3.2, $\operatorname{Cu}(A)$ has $m$-comparison, hence the result follows from Lemma 3.6.
-∎
+Let $A$ be a $C^{*}$-algebra with nuclear dimension at most $m$ for some $m\in{\mathbb{N}}$. Then $A$ has $m$-comparison by [Rob11, Theorem 1.3], hence the result follows from Proposition 3.7. ∎
 4. Variants of divisibility
-We first recall the definition of $n$-almost divisibility for $\mathrm{Cu}$-semigroups, which appeared in [RT17, Section 2.3].
-Other related notions of divisibility will be also useful in our analysis.
-These have sprung into prominence due to their connection with important structural properties of $C^{*}$-algebras.
+In this section we introduce the concept of functional divisibility and show it is a natural generalization of $n$-almost divisibility (as introduced in [Win12, Definition 3.5] for simple, unital $C^{*}$-algebras, and later refined in [RT17, Section 2.3] for general $C^{*}$-algebras).
+We prove that functional divisibility is automatic for nowhere scattered $C^{*}$-algebras of stable rank one, and also for simple, unital, non-elementary $C^{*}$-algebras with a unique normalized quasitrace.
 Definition 4.1.
-Let $k\in{\mathbb{N}}$, and $n\in{\mathbb{N}}\cup\{\infty\}$.
-We say that an element $x$ in a $\mathrm{Cu}$-semigroup $S$ is $(k,n)$-divisible if, for each $x^{\prime}\in S$ with $x^{\prime}\ll x$, there exists $y\in S$ such that
+Let $n\in{\mathbb{N}}$. A $\mathrm{Cu}$-semigroup $S$ is $n$-almost divisible if for every $x^{\prime}\ll x$ in $S$ and every $k\in{\mathbb{N}}$ with $k\geq 1$, there exists $y\in S$ such that
 $ky\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)(n+1)y.$
-If all elements are $(k,n)$-divisible, when $n=\infty$, then we say that $S$ is $(k,\omega)$-divisible.
-If all elements are $(k,n)$-divisible for all $k$, we then say that $S$ is $n$-almost divisible. In case $S$ is 00-almost divisible, then we just say that $S$ is almost divisible.
-A $C^{*}$-algebra is $n$-almost divisible if its Cuntz semigroup is.
-Remarks 4.2.
-(1) It was proved in [TV24, Theorem 9.1] that, if $A$ is a nowhere scattered $C^{*}$-algebra of real rank zero, then $A$ is almost divisible.
-In fact, its Cuntz semigroup $\operatorname{Cu}(A)$ is weakly divisible: for any $x\in\operatorname{Cu}(A)$, there are $y,z\in\operatorname{Cu}(A)$ such that $x=2y+3z$; see also [APGPSM10, Section 6].
-In the simple case, and for the Murray-von Neumann semigroup, this was partly obtained in [PR04, Proposition 5.3].
-(2) If a $\mathrm{Cu}$-semigroup $S$ is almost unperforated, then it is almost divisible precisely when, given $x\in S$ and $k$, there is $y\in S$ such that $ky\leq x\leq(k+1)y$.
-This was proved in [APT18, Proposition 7.3.7].
-(3) The concept of $(k,\omega)$-divisibility is connected to the Global Glimm Property which, in the context of this paper, will be analysed with more detail in Section 6.
-It was proved in [TV23, Theorem 3.6] that a $C^{*}$-algebra has the Global Glimm Property precisely when its Cuntz semigroup is $(2,\omega)$-divisible.
-Inspired by the results in [RT17, Section 6], we introduce the following more general divisibility conditions in terms of comparison by functionals.
+We say that a $C^{*}$-algebra is $n$-almost divisible if its Cuntz semigroup is.
+A $\mathrm{Cu}$-semigroup or $C^{*}$-algebra is said to be almost divisible if it is 00-almost divisible.
+In other words, if for all $x^{\prime}\ll x$ in $S$ and $k\in{\mathbb{N}}$ with $k\geq 1$ there exists $y\in S$ such that $ky\leq x$, and $x^{\prime}\leq(k+1)y$.
+4.2Connections with earlier divisibility notions.
+In parts of the literature, ‘almost divisible’ is defined as the slightly stronger condition that for every $x\in S$ and $k\in{\mathbb{N}}$ with $k\geq 1$ there exists $y\in S$ such that
+$ky\leq x\leq(k+1)y.$
+This version of ‘almost divisibility’ implies the one from Definition 4.1, and for almost unperforated $\mathrm{Cu}$-semigroups both versions agree by [APT18, Proposition 7.3.7].
+The advantage of the version from Definition 4.1 is that it enjoys better permanence properties; see [APT18, Remark 7.3.5].
+Almost divisibility is automatic in certain cases.
+For example, it was shown in [TV24a, Theorem 9.1] that every nowhere scattered $C^{*}$-algebra $A$ of real rank zero is almost divisible.
+In fact, $\operatorname{Cu}(A)$ is even ‘weakly divisible’ in the sense that for every $x\in\operatorname{Cu}(A)$, there are $y,z\in\operatorname{Cu}(A)$ such that $x=2y+3z$; see also [APGPSM10, Section 6].
+In the simple case, and for the Murray-von Neumann semigroup, this was obtained in [PR04, Proposition 5.3].
+Inspired by the results in [RT17, Section 6] and also in [Win12], we introduce the following more general divisibility condition in terms of comparison by functionals.
 Definition 4.3.
-We say that an element $x$ in a $\mathrm{Cu}$-semigroup $S$ is
-  • (i)
-    functionally almost divisible, or that $x$ has (FAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, every $k\in{\mathbb{N}}_{>0}$, and every $\varepsilon>0$, there exists $y\in S$ such that
+We say that a $\mathrm{Cu}$-semigroup $S$ is functionally divisible if the following conditions are satisfied:
+  • (L)
+    For every $x^{\prime}\ll x$ in $S$, every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\varepsilon\in(0,1)$, there exists $y\in S$ such that
+    $ky\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq k\widehat{y}.$
+  • (U)
+    For every $x^{\prime}\ll x$ in $S$, every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\varepsilon\in(0,1)$, there exists $z\in S$ such that
+    $z\leq x,\quad(1-\varepsilon)k\widehat{z}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq kz.$
+We say that a $C^{*}$-algebra $A$ is functionally divisible if $\operatorname{Cu}(A)$ is.
+Remark 4.4.
+Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), and let $x\in S$.
+While Definition 4.1 captures notions of divisibility of $x$ in $S$, one may also consider the divisibility of $\widehat{x}$ in $\operatorname{Lsc}(F(S))$.
+The straightforward concept would be to require that for every $x\in S$ and every $k\in{\mathbb{N}}$ with $k\geq 1$, there exists $y\in S$ such that $k\widehat{y}=\widehat{x}$.
+Taking also the way-below relation into account in order to obtain better permanence properties, one is led to the following notion:
+  • (D)
+    For every $x\in S$, every $f\in\operatorname{Lsc}(F(S))$ with $f\ll\widehat{x}$, and every $k\in{\mathbb{N}}$ with $k\geq 1$, there exists $y\in S$ such that
+    $f\leq k\widehat{y}\leq\widehat{x}.$
+Using Proposition 2.3(i), we see that a function $f\in\operatorname{Lsc}(F(S))$ satisfies $f\ll\widehat{x}$ if and only if there exists $x^{\prime}\in S$ with $x^{\prime}\ll x$ and $\varepsilon>0$ such that $f\leq(1-\varepsilon)\widehat{x^{\prime}}$.
+Thus, condition (D) holds if and only if for every $x^{\prime}\ll x$ in $S$, every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\varepsilon>0$, there exists $y\in S$ such that
+$(1-\varepsilon)\widehat{x^{\prime}}\leq k\widehat{y}\leq\widehat{x}.$
+Conditions (L) and (U) in Definition 4.3 are natural strengthenings of this, where one of the inequalities on functionals is upgraded to an inequality in $S$.
+The ‘L’ stands for ‘lower’, since in this condition the element $ky$ is smaller than $x$, while the ‘U’ stands for ‘upper’ since here $ky$ is larger than $x^{\prime}$.
+The next result shows that it suffices to consider $k=2$ in order to verify condition (L) in Definition 4.3.
+Proposition 4.5.
+Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5). Then the following statements are equivalent:
+  1. (i)
+    $S$ satisfies condition (L) in Definition 4.3.
+  2. (ii)
+    For every $x^{\prime}\ll x$ in $S$ and $k\in{\mathbb{N}}$, there exists $y\in S$ such that
+    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+1\right)\widehat{y}.$
+  3. (iii)
+    For every $x^{\prime}\ll x$ in $S$ and every $\varepsilon\in(0,1)$ there exists $y\in S$ such that
+    $2y\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq 2\widehat{y}.$
+Proof.
+Assuming (i), let us verify (ii).
+For $k=0$, we use $y=x$.
+For $k\geq 1$, we apply (i) with $\varepsilon>0$ small enough such that $(1-\varepsilon)^{-1}k\leq k+1$.
+Similarly, to see that (ii) implies (iii), for a given $\varepsilon\in(0,1)$ apply (ii) with $2k$ where $k$ is large enough such that $1-\varepsilon\leq(1+\tfrac{1}{2k})^{-1}$.
+Next, assuming that (iii) holds, let us verify (i). First, by induction over $m$, we verify that the following statement holds:
+[…]
+    For every $x^{\prime}\ll x$ in $S$ and $\varepsilon\in(0,1)$ there exists $y\in S$ such that
+    $2^{m}y\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq 2^{m}\widehat{y}.$
+[…]
+$x^{\prime}\ll x^{\prime\prime}\ll x,\,\,\,\text{ and }\,\,\,(1-\delta)^{3}=1-\varepsilon.$
+[…]
+$2^{m}w\leq x,\,\,\,\text{ and }\,\,\,(1-\delta)\widehat{x^{\prime\prime}}\leq 2^{m}\widehat{w}.$
+By Proposition 2.3(i), we have $(1-\delta)^{2}\widehat{x^{\prime}}\ll(1-\delta)\widehat{x^{\prime\prime}}$ in $\operatorname{Lsc}(F(S))$, hence also $(1-\delta)^{2}\widehat{x^{\prime}}\ll 2^{m}\widehat{w}$.
+This allows us to choose $w^{\prime}\in S$ such that
+$w^{\prime}\ll w,\,\,\,\text{ and }\,\,\,(1-\delta)^{2}\widehat{x^{\prime}}\leq 2^{m}\widehat{w^{\prime}}.$
+Next, applying (iii) for $w^{\prime}\ll w$ and $\delta$, we obtain $y\in S$ such that
+$2y\leq w,\,\,\,\text{ and }\,\,\,(1-\delta)\widehat{w^{\prime}}\leq 2\widehat{y}.$
+Then
+$2^{m+1}y\leq 2^{m}w\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}=(1-\delta)^{3}\widehat{x^{\prime}}\leq(1-\delta)2^{m}\widehat{w^{\prime}}\leq 2^{m+1}\widehat{y},$
+[…]
+Now, to prove (i), let $x^{\prime}\ll x$ in $S$, let $k\in{\mathbb{N}}$ with $k\geq 1$, and let $\varepsilon\in(0,1)$. Pick $x^{\prime\prime}\in S$ and $\delta>0$ such that
+$x^{\prime}\ll x^{\prime\prime}\ll x,\,\,\,\text{ and }\,\,\,(1-\delta)^{2}=1-\varepsilon.$
+Using that dyadic rationals are dense in $\mathbb{R}$, choose $n,m\geq 1$ such that
+$(1-\delta)\frac{1}{k}\leq\frac{n}{2^{m}}\leq\frac{1}{k}.$
+Then
+$(1-\delta)2^{m}\leq kn\leq 2^{m}.$
+[…]
+$2^{m}w\leq x,\,\,\,\text{ and }\,\,\,(1-\delta)\widehat{x^{\prime}}\leq 2^{m}\widehat{w}.$
+Set $y:=nw$. Then
+$ky=k(nw)\leq 2^{m}w\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}=(1-\delta)^{2}\widehat{x^{\prime}}\leq(1-\delta)2^{m}\widehat{w}\leq kn\widehat{w}=k\widehat{y},$
+which shows that $y$ has the desired properties. ∎
+Remark 4.6.
+Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5).
+It is not clear if there exists a characterization of condition (U) in Definition 4.3 analogous to Proposition 4.5.
+The reason is that condition (U) has the additional assumption that the dividing element is dominated by $x$, which is automatic in condition (L).
+On the other hand, for $m\geq 1$ let us consider the following condition:
+  • ($U_{2^{m}}$)
+    For every $x^{\prime}\ll x$ in $S$ and every $\varepsilon\in(0,1)$, there exists $y\in S$ such that
+    $y\leq x,\quad(1-\varepsilon)2^{m}\widehat{y}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq 2^{m}y.$
+An argument as in the proof of Proposition 4.5 shows that if $S$ satisfies ($U_{2}$), then it satisfies ($U_{2^{m}}$) for all $m$. (See also Remark 5.6.)
+To show that $n$-almost divisible Cuntz semigroups are functionally divisible (Proposition 4.9), we first need some preparatory technical results.
+We will use the concept of $(2,\omega)$-divisibility (see also Section 6).
+Given $k\in{\mathbb{N}}$ with $k\geq 2$, a $\mathrm{Cu}$-semigroup $S$ is said to be $(k,\omega)$-divisible if for all $x^{\prime}\ll x$ in $S$ there exists $y\in S$ such that
+$ky\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq\infty y.$
+If $S$ is $(k,\omega)$-divisible for some $k\geq 2$, then it is $(2,\omega)$-divisible.
+The converse also holds by [TV23, Lemma 3.4].
+Thus, a $\mathrm{Cu}$-semigroup is $(2,\omega)$-divisible if and only if it is $(k,\omega)$-divisible for all $k\geq 2$.
+Condition (L’) in the result below was considered in [RT17, Theorem 6.1].
+Lemma 4.7.
+Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5)-(O7). Then $S$ satisfies condition (L) in Definition 4.3 if and only if $S$ is $(2,\omega)$-divisible and satisfies
+  1. (L’)
+    For every $x^{\prime}\ll x$ in $S$, every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\varepsilon>0$, there exists $y\in S$ such that
     $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq k\widehat{y}+\varepsilon\widehat{x}.$
-  • (ii)
-    lower functionally almost divisible, or that $x$ has (LFAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
-    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+1\right)\widehat{y}.$
-  • (iii)
-    upper functionally almost divisible, or that $x$ has (UFAD), if for every $x^{\prime}\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ and $\nu>0$ such that
-    $y\leq x,\quad(k+\nu)\widehat{y}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)y.$
-A $\mathrm{Cu}$-semigroup $S$ is said to be (upper, lower) functionally almost divisible if every element in $S$ is (upper, lower) functionally almost divisible.
-We begin our study of functional divisibility by providing a number of technical results relating each of the notions above.
-These results will be applied to $C^{*}$-algebras with $n$-almost divisibility at the end of the section, where we will show that any $n$-almost divisible $C^{*}$-algebra satisfies all three divisibility notions; see Theorem 4.7.
-Lemma 4.4.
-Let $S$ be a $\mathrm{Cu}$-semigroup.
-  1. (i)
-    $S$ is functionally almost divisible if and only if for every $x^{\prime},x\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
-    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq k\widehat{y}+\tfrac{1}{k}\widehat{x}.$
-  2. (ii)
-    $S$ is lower functionally almost divisible if and only if for every $x^{\prime},x\in S$ satisfying $x^{\prime}\ll x$, and every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
-    $ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq(k+\tfrac{1}{k})\widehat{y}.$
 Proof.
-We will only prove (i) since the proof of (ii) is similar.
-Moreover the forward implications are clear.
-To prove the backwards implication, let $x^{\prime},x\in S$ satisfy $x^{\prime}\ll x$, let $k\in{\mathbb{N}}_{>0}$, and let $\varepsilon>0$.
-Choose $n\in{\mathbb{N}}_{>0}$ such that $\tfrac{1}{kn}\leq\varepsilon$.
-Applying the assumption for $x^{\prime},x$ and $kn$, we obtain $z\in S$ such that
-$knz\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq kn\widehat{z}+\frac{1}{kn}\widehat{x}.$
-Set $y:=nz$. Then
-$ky=knz\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq kn\widehat{z}+\frac{1}{kn}\widehat{x}=k\widehat{y}+\frac{1}{kn}\widehat{x}\leq k\widehat{y}+\varepsilon\widehat{x},$
-as desired. ∎
-Lemma 4.5.
-Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5) and (O7). Assume that $S$ is $(2,\omega)$-divisible and is functionally almost divisible. Then $S$ is lower functionally almost divisible.
-Proof.
-Let $k\in{\mathbb{N}}$, and let $x^{\prime},x\in S$ be such that $x^{\prime}\ll x$. It suffices to prove that there exists $y\in S$ such that
-$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+\frac{1}{k}\right)\widehat{y}.$
-For $k=1$ use $y:=x$. Thus, we may from now on assume that $k\geq 2$.
-Using the methods from [TV21, Section 5], we find a countably based, $(2,\omega)$-divisible $\mathrm{Cu}$-semigroup $H$ that contains $x^{\prime},x$ and such that $H$ satisfies (O5), (O7), and (FAD).
-Choose $x^{\prime\prime}\in H$ and $n\in{\mathbb{N}}_{>0}$ such that
-$x^{\prime}\ll x^{\prime\prime}\ll x,\quad k+1\leq n,\,\,\,\text{ and }\,\,\,\frac{k(k+1)+n}{n}\leq 1+\frac{1}{k^{2}}.$
+To show the forward implication, assume that $S$ satisfies
+  • (L)
+    For every $x^{\prime}\ll x$ in $S$, every $k\in{\mathbb{N}}\setminus\{0\}$, and every $\varepsilon\in(0,1)$, there exists $y\in S$ such that
+    $ky\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq k\widehat{y}.$
+To verify that $S$ is $(2,\omega)$-divisible, let $x^{\prime}\ll x$ in $S$. Applying (L) with the given $x^{\prime},x$, and also $k=2$ and $\varepsilon=\tfrac{1}{2}$, we obtain $y\in S$ such that
+$2y\leq x,\,\,\,\text{ and }\,\,\,\frac{1}{2}\widehat{x^{\prime}}\leq 2\widehat{y}.$
+The latter condition implies that $\widehat{x^{\prime}}\leq 4\widehat{y}$, and hence $x^{\prime}\leq\infty y$ (using, for example, [APRT22, Lemma 6.6]). This shows that $S$ is $(2,\omega)$-divisible.
+To verify (L’), let $x^{\prime}\ll x$ in $S$, let $k\in{\mathbb{N}}$ with $k\geq 1$, and $\varepsilon>0$.
+Pick $\delta\in(0,1)$ such that $(1-\delta)^{-1}=1+\varepsilon$.
+Applying (L) with the given $x^{\prime},x$ and $k$, and also $\delta$, we obtain $y\in S$ such that
+$ky\leq x,\,\,\,\text{ and }\,\,\,(1-\delta)\widehat{x^{\prime}}\leq k\widehat{y}.$
+Using that $ky\leq x$, we get
+$\widehat{x^{\prime}}\leq(1-\delta)^{-1}k\widehat{y}=(1+\varepsilon)k\widehat{y}\leq k\widehat{y}+\varepsilon\widehat{x},$
+as desired.
+To show the backward implication, assume that $S$ is $(2,\omega)$-divisible and satisfies (L’).
+We verify condition (ii) of Proposition 4.5.
+Let $x^{\prime}\ll x$ in $S$, and let $k\in{\mathbb{N}}$.
+We need to find $y\in S$ such that
+$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq(k+1)\widehat{y}.$
+For $k=0$ and $k=1$ use $y:=x$. Thus, from now on we may assume that $k\geq 2$.
+Using the methods from [TV21, Section 5], we find a countably based, $(2,\omega)$-divisible sub-$\mathrm{Cu}$-semigroup $H\subseteq S$ that satisfies (O5)-(O7) and contains $x^{\prime}$, $x$.
+Choose $x^{\prime\prime}\in H$ and $n\in{\mathbb{N}}$ with $n\geq 1$ such that
+$x^{\prime}\ll x^{\prime\prime}\ll x,\quad k+1\leq n,\,\,\,\text{ and }\,\,\,\frac{k(k+1)+n}{n}\leq 1+\frac{1}{k}.$
 Since $H$ is $(2,\omega)$-divisible, it is also $(n,\omega)$-divisible by the arguments in [TV23, Lemma 3.4]. Applied to $x^{\prime\prime}\ll x$ and $n$, we obtain $c\in H$ such that
@@
 $nx^{\prime\prime}\leq nx\leq nkc+nd\leq kx+nd\leq(k(k+1)+n)d.$
-Since $H$ is countably based and satisfies (O7), we may apply Theorem 2.2.
-Hence, for every $s\in S$ the infimum $s\wedge\infty c^{\prime}$ exists, and the map $s\mapsto s\wedge\infty c^{\prime}$ is additive; Set $e:=d\wedge\infty c^{\prime}$.
+Since $H$ is countably based and satisfies (O5)-(O7), we may apply Theorem 2.2.
+Hence, for every $s\in H$ the infimum $s\wedge\infty c^{\prime}$ exists (in $H$), and the map $s\mapsto s\wedge\infty c^{\prime}$ is additive.
+Set $e:=d\wedge\infty c^{\prime}$.
 Since $x^{\prime\prime}\ll\infty c^{\prime}$, we have
@@
 $nx^{\prime}\leq(k(k+1)+n)e^{\prime},\,\,\,\text{ and }\,\,\,e^{\prime}\ll e.$
-Let us further choose $e^{\prime\prime}$ such that $e^{\prime}\ll e^{\prime\prime}\ll e$.
-There exists $N\in{\mathbb{N}}$ such that $e^{\prime\prime}\leq Nc^{\prime}$. Choose $\varepsilon>0$ small enough such that
-$\left(1+\frac{1}{k^{2}}\right)\varepsilon\leq\frac{1}{N}.$
-Applying (FAD) to $e^{\prime}\ll e^{\prime\prime}$, $k$ and $\varepsilon$, we obtain $f\in H$ such that
-$kf\leq e^{\prime\prime},\,\,\,\text{ and }\,\,\,\widehat{e^{\prime}}\leq\widehat{kf}+\varepsilon\widehat{e^{\prime\prime}}.$
+Let us further choose $e^{\prime\prime}\in H$ such that $e^{\prime}\ll e^{\prime\prime}\ll e$.
+Since $e\leq\infty c^{\prime}$ and $e^{\prime}\ll e$, we may pick $N\in{\mathbb{N}}$ with $N\geq 1$ such that $e^{\prime\prime}\leq Nc^{\prime}$. Then choose $\varepsilon>0$ small enough such that
+$\left(1+\frac{1}{k}\right)\varepsilon\leq\frac{1}{N}.$
+Now, working again in $S$, and applying (L’) to $e^{\prime}\ll e^{\prime\prime}$, $k$ and $\varepsilon$, we obtain $f\in S$ such that
+$kf\leq e^{\prime\prime},\,\,\,\text{ and }\,\,\,\widehat{e^{\prime}}\leq k\widehat{f}+\varepsilon\widehat{e^{\prime\prime}}.$
 Set $y:=c^{\prime}+f$. We have $kf\leq e\leq d$ and therefore
 $ky=kc^{\prime}+kf\leq kc^{\prime}+d\leq x.$
-To verify that $\widehat{x^{\prime}}\leq(k+1/k)\widehat{y}$, let $\lambda\in F(H)$. Then
-$\begin{split}\lambda(x^{\prime})&\leq\frac{k(k+1)+n}{n}\lambda(e^{\prime})\leq\left(1+\frac{1}{k^{2}}\right)\lambda(e^{\prime})\\ &\leq\left(1+\frac{1}{k^{2}}\right)\left(k\lambda(f)+\varepsilon\lambda(e^{\prime\prime})\right)\leq\left(k+\frac{1}{k}\right)\lambda(f)+\frac{1}{N}\lambda(e^{\prime\prime})\\ &\leq\left(k+\frac{1}{k}\right)\lambda(f)+\lambda(c^{\prime})\leq\left(k+\frac{1}{k}\right)\lambda(y),\end{split}$
+Further, we have
+$\widehat{x^{\prime}}$ $\leq\frac{k(k+1)+n}{n}\widehat{e^{\prime}}\leq\left(1+\frac{1}{k}\right)\widehat{e^{\prime}}\leq\left(1+\frac{1}{k}\right)\left(k\widehat{f}+\varepsilon\widehat{e^{\prime\prime}}\right)$
+$\leq(k+1)\widehat{f}+\frac{1}{N}\widehat{e^{\prime\prime}}\leq(k+1)\widehat{f}+\widehat{c^{\prime}}\leq(k+1)\left(\widehat{f}+\widehat{c^{\prime}}\right)=(k+1)\widehat{y},$
 as desired. ∎
+The next result shows that if a compact element satisfies condition (L’) in Lemma 4.7, then it also satisfies condition (U) in Definition 4.3.
 Lemma 4.8.
-Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), and let $x\in S$ be a compact element.
-If $x$ is functionally almost divisible, then for every $k\in{\mathbb{N}}_{>0}$, there exists $y\in S$ such that
-$y\leq x,\quad(k+\tfrac{1}{2})\widehat{y}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x\leq(k+1)y.$
-In particular, $x$ is upper functionally almost divisible.
+Let $S$ be a $\mathrm{Cu}$-semigroup satisfying (O5), and let $x\in S$ be a compact element such that for every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\varepsilon>0$, there exists $y\in S$ such that
+$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x}\leq k\widehat{y}+\varepsilon\widehat{x}.$
+Then, for every $k\in{\mathbb{N}}$ with $k\geq 1$, and every $\delta>0$, there exists $z\in S$ such that
+$z\leq x,\quad(1-\delta)k\widehat{z}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x\leq kz.$
 Proof.
-Let $k\in{\mathbb{N}}$. Using (FAD) with $k+1$ and $\varepsilon=1/3(2k+1)(k+1)$, we obtain an element $z\in S$ such that
-$(k+1)z\leq x,\,\,\,\text{ and }\,\,\,\widehat{x}\leq(k+1)\widehat{z}+\frac{1}{3(2k+1)(k+1)}\widehat{x}.$
-Find $\delta>0$ such that
-$\frac{k+1}{1-\delta}\leq(k+1)+\frac{1}{2(2k+1)},\,\,\,\text{ and }\,\,\,\frac{\varepsilon}{(1-\delta)}\leq\frac{1}{2(2k+1)(k+1)}.$
+Given $k\in{\mathbb{N}}$ with $k\geq 1$ and $\delta>0$, pick $\varepsilon>0$ such that
+$1+2(k-1)\varepsilon\leq(1-\delta)^{-1}.$
+Applying the assumption for $k$ and $\varepsilon$, we obtain $y\in S$ such that
+$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x}\leq k\widehat{y}+\varepsilon\widehat{x}.$
 Now, using that $x\ll x$, it follows from Proposition 2.3(i) that
-$(1-\delta)\widehat{x}\ll\widehat{x}\leq(k+1)\widehat{z}+\varepsilon\widehat{x}.$
-This allows us to choose $z^{\prime}\ll z$ such that
-(1) $\widehat{x}\leq\frac{k+1}{(1-\delta)}\widehat{z^{\prime}}+\frac{\varepsilon}{(1-\delta)}\widehat{x}\leq\left((k+1)+\frac{1}{2(2k+1)}\right)\widehat{z^{\prime}}+\frac{1}{2(2k+1)(k+1)}\widehat{x}$
-Using Lemma 2.1, we can now find $y\in S$ such that
-$kz^{\prime}+y\leq x\leq(k+1)y.$
-Take $\lambda\in F(S)$. If $\lambda(z^{\prime})=\infty$, we have $\lambda(y)=\lambda(x)=\infty$ and so, in particular, $(k+\tfrac{1}{2})\lambda(y)\leq\lambda(x)$.
-Thus, assume that $\lambda(z^{\prime})<\infty$. Since $k\lambda(y)\leq k(\lambda(x)-k\lambda(z^{\prime}))$, using (1) at the first step, and $(k+1)z^{\prime}\leq x$ at the last step, one has:
-$\begin{split}(k+\tfrac{1}{2})\lambda(y)&\leq(k+\tfrac{1}{2})\left(\left((k+1)+\tfrac{1}{2(2k+1)}\right)\lambda(z^{\prime})+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)-k\lambda(z^{\prime})\right)\\ &=(k+\tfrac{1}{2})\left(\left(1+\tfrac{1}{2(2k+1)}\right)\lambda(z^{\prime})+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)\right)\\ &\leq(k+\tfrac{1}{2})\left(\tfrac{1}{k+1}\left(1+\tfrac{1}{2(2k+1)}\right)\lambda(x)+\tfrac{1}{2(2k+1)(k+1)}\lambda(x)\right)=\lambda(x),\end{split}$
-as desired.
-The last assertion is clear. ∎
-Theorem 4.7.
-Let $A$ be a $C^{*}$-algebra. Assume that $A$ is $n$-almost divisible for some $n\in{\mathbb{N}}$. Then $\operatorname{Cu}(A)$ is (upper, lower) functionally almost divisible.
+$(1-\varepsilon)\widehat{x}\ll\widehat{x}\leq k\widehat{y}+\varepsilon\widehat{x}.$
+This allows us to choose $y^{\prime}\in S$ such that $y^{\prime}\ll y$ and
+$(1-\varepsilon)\widehat{x}\leq k\widehat{y^{\prime}}+\varepsilon\widehat{x}.$
+Applying Lemma 2.1 for $ky\leq x$ and $y^{\prime}\ll y$, we get $z\in S$ such that
+$(k-1)y^{\prime}+z\leq x\leq kz.$
+We now show that $z$ has the claimed properties.
+It is clear that $z\leq x$ and $x\leq kz$.
+It remains to verify that $(1-\delta)k\lambda(z)\leq\lambda(x)$ for every $\lambda\in F(S)$.
+This is clear if $\lambda(x)=\infty$.
+Let $\lambda\in F(S)$ with $\lambda(x)<\infty$.
+From the fact that $(1-\varepsilon)\widehat{x}\leq k\widehat{y^{\prime}}+\varepsilon\widehat{x}$, as shown above, we deduce that
+$(1-2\varepsilon)\lambda(x)\leq k\lambda(y^{\prime}).$
+Since $ky\leq x$, it follows that $\lambda(y^{\prime})<\infty$ and, consequently, using our choice of $\varepsilon$ in the last step,
+$k\lambda(z)$ $\leq k\lambda(x)-k(k-1)\lambda(y^{\prime})\leq k\lambda(x)-(1-2\varepsilon)(k-1)\lambda(x)$
+$=(1+2(k-1)\varepsilon)\lambda(x)\leq(1-\delta)^{-1}\lambda(x),$
+as desired. ∎
+Proposition 4.9.
+If a $C^{*}$-algebra is $n$-almost divisible for some $n\in{\mathbb{N}}$, then it is functionally divisible.
 Proof.
-Assume that $A$ is $n$-almost divisible.
-Without loss of generality, we may assume $A$ to be stable.
-Further, note that we only need to prove that $\operatorname{Cu}(A)$ has (FAD) and (UFAD).
-Indeed, any $n$-almost divisible $C^{*}$-algebra has a $(2,\omega)$-divisible Cuntz semigroup.
-Thus, it follows from Lemma 4.5 that $\operatorname{Cu}(A)$ has (LFAD) whenever it has (FAD).
-Let $\mathcal{U}$ be a free ultrafilter on ${\mathbb{N}}$, and let $a\in A_{+}$.
+Let $n\in{\mathbb{N}}$, and let $A$ be a $C^{*}$-algebra that is $n$-almost divisible.
+It is easy to deduce that $\operatorname{Cu}(A)$ is $(2,\omega)$-divisible.
+Therefore, by Lemma 4.7, to show that $\operatorname{Cu}(A)$ is functionally divisible it suffices to verify condition (L’) in Lemma 4.7 and condition (U) in Definition 4.3.
+Let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, let $k\in{\mathbb{N}}$ with $k\geq 1$, and let $\varepsilon>0$.
+We may assume that $\varepsilon=\tfrac{1}{N}$ for some integer $N\geq 1$, and we need to find $y\in\operatorname{Cu}(A)$ such that
+$ky\leq x,\,\,\,\text{ and }\,\,\,N\widehat{x^{\prime}}\leq Nk\widehat{y}+\widehat{x},$
+and $z\in\operatorname{Cu}(A)$ such that
+$z\leq x,\quad(N-1)k\widehat{z}\leq N\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq kz.$
+Without loss of generality, we may assume that $A$ is stable, which allows us to pick $a\in A_{+}$ with $x=[a]$.
+Let $\mathcal{U}$ be a free ultrafilter on ${\mathbb{N}}$.
+Denote by $A_{\mathcal{U}}$ the free ultrapower of $A$.
 Set $C=\{a\}^{\prime}\cap A_{\mathcal{U}}$ and $I=\{a\}^{\perp}\cap A_{\mathcal{U}}$.
-Then, it follows from [RT17, Theorem 6.1] that the class of the unit $[1]$ in $\operatorname{Cu}(C/I)$ has (FAD).
-Therefore, $[1]$ has (UFAD) by Lemma 4.6.
-Now choose $k\in{\mathbb{N}}_{>0}$. Applying (FAD) to $[1]\ll[1]$ with $\varepsilon=\tfrac{1}{k+1}$, we obtain an element $e\in C/I$ such that
-$k[e]\leq[1],\,\,\,\text{ and }\,\,\,\widehat{[1]}\leq k\widehat{[e]}+\tfrac{1}{k+1}\widehat{[1]}.$
-Setting $\delta=\tfrac{1}{k(k+1)}$, we have $(1+\delta)\widehat{k[1]}\leq k^{2}\widehat{[e]}+\widehat{[1]}$.
-Now we apply Proposition 2.3(ii) to find a positive integer $s$ such that $(s+1)k[1]\leq s(k^{2}[e]+[1])$.
-Separately, we apply (UFAD) to $[1]\ll[1]$ and we obtain an element $\tilde{f}\in C/I$ such that
-$[\tilde{f}]\leq[1],\quad(k+\nu)\widehat{[\tilde{f}]}\leq\widehat{[1]},\,\,\,\text{ and }\,\,\,[1]\leq(k+1)[\tilde{f}],$
-for some $10<\nu<$ 1.
-Since $[1]\ll[1]$, we can find $f\in C/I$ with $[f]\ll[\tilde{f}]$ such that $[1]\leq(k+1)[f]$.
-Again, applying Proposition 2.3(ii), we find a positive integer $t$ such that $(t+1)k[f]\leq t[1]$.
-Let $\bar{e},\bar{f}\in C$ be lifts of $e,f$ respectively. Then, there exists $z\in\operatorname{Cu}(I)$ such that
-$k[\bar{e}]\leq[1]+z,\quad(s+1)k[1]\leq s(k^{2}[\bar{e}]+[1])+z,$
+Then, it follows from [RT17, Theorem 6.1] that the class of the unit $[1]$ in $\operatorname{Cu}(C/I)$ satisfies condition (L’) in Lemma 4.7, that is, for every $l\in{\mathbb{N}}$ with $l\geq 1$, and every $\delta>0$, there exists $v\in\operatorname{Cu}(C/I)$ such that
+$lv\leq[1],\,\,\,\text{ and }\,\,\,\widehat{[1]}\leq l\widehat{v}+\delta\widehat{[1]}.$
+Applied for $l=k$ and $\delta=\tfrac{1}{2N}$, we obtain $v\in\operatorname{Cu}(C/I)$ such that
+$kv\leq[1],\,\,\,\text{ and }\,\,\,\widehat{[1]}\leq k\widehat{v}+\frac{1}{2N}\widehat{[1]}.$
+Adding $\tfrac{1}{2N}\widehat{[1]}$ and multiplying by $N$, we get
+$(1+\frac{1}{2N})N\widehat{[1]}\leq Nk\widehat{v}+\widehat{[1]}.$
+Now we apply Proposition 2.3(ii) to find a positive integer $s$ such that
+$sN[1]\leq(s+1)N[1]\leq s(Nkv+[1]).$
+Since $sN[1]$ is compact, we can pick $v^{\prime}\in\operatorname{Cu}(C/I)$ such that
+$v^{\prime}\ll v,\,\,\,\text{ and }\,\,\,sN[1]\leq s(Nkv^{\prime}+[1]).$
+Using also that $v\leq[1]$ (since $kv\leq[1]$), we can apply [RR13, Lemma 2.3 (ii)] to find $e\in(C/I)_{+}$ such that $v^{\prime}\leq[e]\leq v$, and so
+$k[e]\leq[1],\,\,\,\text{ and }\,\,\,sN[1]\leq s(Nk[e]+[1]).$
+Separately, and using again that $[1]\in\operatorname{Cu}(C/I)$ satisfies condition (L’), we apply Lemma 4.8 to the compact element $[1]$ in $\operatorname{Cu}(C/I)$ and for the given $k$ and $\tfrac{1}{2N}$ to obtain $w\in\operatorname{Cu}(C/I)$ such that
+$w\leq[1],\quad(1-\tfrac{1}{2N})k\widehat{w}\leq\widehat{[1]},\,\,\,\text{ and }\,\,\,[1]\leq kw.$
+Multiplying by $N$, we get
+$(N-\tfrac{1}{2})k\widehat{w}\leq N\widehat{[1]}.$
+Using that $[1]$ is compact and $[1]\leq kw$, we find $w^{\prime\prime},w^{\prime}\in\operatorname{Cu}(C/I)$ such that
+$w^{\prime\prime}\ll w^{\prime}\ll w,\,\,\,\text{ and }\,\,\,[1]\leq kw^{\prime\prime}.$
+First, applying Proposition 2.3(ii), we find a positive integer $t$ such that
+$t(N-1)kw^{\prime}\leq(t+1)(N-1)kw^{\prime}\leq tN[1].$
+Then, arguing as above, we find $f\in(C/I)_{+}$ with $w^{\prime\prime}\leq[f]\leq w^{\prime}$, and thus
+$[f]\leq[1],\quad t(N-1)k[f]\leq tN[1],\,\,\,\text{ and }\,\,\,[1]\leq k[f].$
+Let $\bar{e},\bar{f}\in C_{+}$ be lifts of $e,f$ respectively. Then, there exists $z\in\operatorname{Cu}(I)$ such that
+$k[\bar{e}]\leq[1]+z,\quad sN[1]\leq s(Nk[\bar{e}]+[1])+z,$
 and
-$[\bar{f}]\leq[1]+z,\quad(t+1)k[\bar{f}]\leq t[1]+z,\quad[1]\leq(k+1)[\bar{f}]+z$
+$[\bar{f}]\leq[1]+z,\quad t(N-1)k[\bar{f}]\leq tN[1]+z,\quad[1]\leq k[\bar{f}]+z$
 in $\operatorname{Cu}(C)$.
@@
 Using that every element in $C$ commutes with $a$ (in particular, those elements applying the Cuntz subequivalences above) and that every element in $I$ is orthogonal to $a$, one obtains
-$k[b]\leq[a],\quad(s+1)k[a]\leq s(k^{2}[b]+[a]),$
+$k[b]\leq[a],\quad sN[a]\leq s(Nk[b]+[a]),$
 and
-$[c]\leq[a],\quad(t+1)k[c]\leq t[a],\quad[a]\leq(k+1)[c]$
+$[c]\leq[a],\quad t(N-1)k[c]\leq tN[a],\quad[a]\leq k[c]$
 in $\operatorname{Cu}(A_{\mathcal{U}})$.
-This shows that $[a]\in\operatorname{Cu}(A_{\mathcal{U}})$ has (FAD) and (UFAD).
-Now, using the fact that functional comparison can be restated in terms of comparison in the semigroup ([Rob13, Proposition 2.2.2]), a standard model theoretic argument shows that $[a]\in\operatorname{Cu}(A)$ also has (FAD) and (UFAD); see [FHL+21, Proposition 8.1.6] and also [APRT24, Theorem 10.5].
+A finite collection of Cuntz subequivalences in $A_{\mathcal{U}}$ can be ‘lifted’ to $A$ simultaneously up to a cut-down (see, for example, the proof of [APRT24, Theorem 10.5]).
+Thus, using that $x^{\prime}\ll x=[a]$, we find $y,z\in\operatorname{Cu}(A)$ such that
+$ky\leq x,\quad sNx^{\prime}\leq s(Nky+x),$
+and
+$z\leq x,\quad t(N-1)kz\leq tNx,\quad x^{\prime}\leq kz$
+in $\operatorname{Cu}(A)$.
+Passing to functionals, we can remove the multiplication by $s$ and $t$, and we obtain $N\widehat{x^{\prime}}\leq Nk\widehat{y}+\widehat{x}$ and $(N-1)k\widehat{z}\leq N\widehat{x}$, which shows that $y$ and $z$ have the desired properties.
 ∎
+It was proved in [APRT22, Theorem 3.8] that, if $A$ is a separable $C^{*}$-algebra of stable rank one, then $\operatorname{Cu}(A)$ is an inf-semilattice ordered semigroup.
+That is, for each $x,y\in\operatorname{Cu}(A)$ their greatest lower bound $x\wedge y$ exists in $\operatorname{Cu}(A)$ and addition is distributive over the meet operation:
+$(x\wedge y)+z=(x+z)\wedge(y+z)$
+for all $x,y,z\in\operatorname{Cu}(A)$. This will be used below.
+Proposition 4.10.
+Every nowhere scattered $C^{*}$-algebra with stable rank one is functionally divisible.
+Proof.
+Let us first assume that every separable, nowhere scattered $C^{*}$-algebra with stable rank one is functionally divisible, and let $A$ be an arbitrary nowhere scattered $C^{*}$-algebra with stable rank one.
+To show that $A$ is functionally divisible, let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, let $k\in{\mathbb{N}}$ with $k\geq 1$ and let $\varepsilon\in(0,1)$ be given.
+We need to find $y,z\in\operatorname{Cu}(A)$ such that
+$ky\leq x,\quad(1-\varepsilon)\widehat{x^{\prime}}\leq k\widehat{y},\quad z\leq x,\quad(1-\varepsilon)k\widehat{z}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq kz.$
+Using that stable rank one and nowhere scatteredness satisfy the Löwenheim-Skolem condition ([TV24a, Proposition 4.11]) and applying [TV21, Proposition 6.1], we can find a separable sub-$C^{*}$-algebra $B\subseteq A$ such that $B$ is nowhere scattered, has stable rank one, and such that the inclusion $B\to A$ induces an order-embedding $\operatorname{Cu}(B)\to\operatorname{Cu}(A)$ whose image contains $x^{\prime}$ and $x$.
+Viewing $x^{\prime}$ and $x$ as elements in $\operatorname{Cu}(B)$, and using that $B$ is functionally divisible, we find $y$ and $z$ with the desired properties in $\operatorname{Cu}(B)$, and thus in $\operatorname{Cu}(A)$.
+(See also [APRT22, Lemma 9.2].)
+By the argument above, we may assume that $A$ is a separable, nowhere scattered $C^{*}$-algebra $A$ with stable rank one, and we need to show that $A$ is functionally divisible.
+We first verify condition (iii) in Proposition 4.5, which then implies that $\operatorname{Cu}(A)$ satisfies condition (L).
+Let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, and let $\varepsilon>0$. We need to find $y\in\operatorname{Cu}(A)$ such that
+$2y\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq 2\widehat{y}.$
+Using the solution to the rank problem (see Theorem 2.4) we can find $u\in\operatorname{Cu}(A)$ such that $\widehat{u}=\tfrac{1}{2}\widehat{x}$.
+Set $v:=u\wedge x$, which exists by the comments preceding this proposition.
+Using [APRT22, Theorem 6.12] at the second step, we have
+$\widehat{v}=\widehat{u\wedge x}=\widehat{u}\wedge\widehat{x}=\frac{1}{2}\widehat{x}.$
+Using that $\operatorname{Cu}(A)$ is $(2,\omega)$-divisible by the results in [APRT22, Section 5] (see also Theorem 7.1 and Proposition 7.3 in [TV23]), we can apply [AVTV23, Theorem 5.10] to find a soft element $w\in\operatorname{Cu}(A)$ such that
+$w\leq v,\,\,\,\text{ and }\,\,\,\widehat{w}=\widehat{v}.$
+By Proposition 2.3(i), we have
+$(1-\varepsilon)\frac{1}{2}\widehat{x^{\prime}}\ll\frac{1}{2}\widehat{x}=\widehat{w},$
+which allows us to choose $w^{\prime\prime},w^{\prime}\in\operatorname{Cu}(A)$ such that
+$w^{\prime\prime}\ll w^{\prime}\ll w,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\frac{1}{2}\widehat{x^{\prime}}\leq\widehat{w^{\prime\prime}}.$
+Since $w$ is soft, there exists by [TV24b, Proposition 4.6] some $t\in\operatorname{Cu}(A)$ such that
+$w^{\prime}+t\ll w,\,\,\,\text{ and }\,\,\,w^{\prime}\ll\infty t.$
+In particular, $w^{\prime}+t\leq v\leq x$.
+Choose $t^{\prime}\in\operatorname{Cu}(A)$ such that
+$t^{\prime}\ll t,\,\,\,\text{ and }\,\,\,w^{\prime}\ll\infty t^{\prime}.$
+Applying (O5) for $w^{\prime}+t\leq x$ and $w^{\prime\prime}\ll w^{\prime}$ and $t^{\prime}\ll t$, we obtain $c\in\operatorname{Cu}(A)$ such that
+$w^{\prime\prime}+c\leq x\leq w^{\prime}+c,\,\,\,\text{ and }\,\,\,t^{\prime}\ll c.$
+Let us show that $\tfrac{1}{2}\widehat{x}\leq\widehat{c}$.
+Since $w^{\prime}\ll\infty t^{\prime}$, there is $m\geq 1$ such that $w^{\prime}\leq mt^{\prime}$, and thus $x\leq w^{\prime}+c\leq mt^{\prime}+t^{\prime}=(m+1)t^{\prime}$.
+Therefore, if $\lambda\in F(\operatorname{Cu}(A))$ satisfies $\lambda(x)=\infty$, then $\lambda(c)=\infty\geq\tfrac{1}{2}\lambda(x)$.
+On the other hand, if $\lambda\in F(\operatorname{Cu}(A))$ satisfies $\lambda(x)<\infty$, then using that $w^{\prime}\leq w$ and $\widehat{w}=\tfrac{1}{2}\widehat{x}$, we get
+$\lambda(c)\geq\lambda(x)-\lambda(w^{\prime})\geq\lambda(x)-\frac{1}{2}\lambda(x)=\frac{1}{2}\lambda(x).$
+Set $y:=w^{\prime\prime}\wedge c$.
+Using that $(1-\varepsilon)\tfrac{1}{2}\widehat{x^{\prime}}\leq\widehat{w^{\prime\prime}}$ and $\tfrac{1}{2}\widehat{x}\leq\widehat{c}$, and using [APRT22, Theorem 6.12] at the second step, we have
+$(1-\varepsilon)\tfrac{1}{2}\widehat{x^{\prime}}\leq\widehat{w^{\prime\prime}}\wedge\widehat{c}=\widehat{w^{\prime\prime}\wedge c}=\widehat{y},$
+and thus
+$2y\leq w^{\prime\prime}+c\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq 2\widehat{y},$
+as desired.
+Next, to verify condition (U) from Definition 4.3, let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, let $k\in{\mathbb{N}}$ with $k\geq 1$, and let $\varepsilon\in(0,1)$.
+We need to find $z\in S$ such that
+$z\leq x,\quad(1-\varepsilon)k\widehat{z}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq kz.$
+Set $s:=(1-\varepsilon)^{-1}\tfrac{1}{k}$.
+If $s\geq 1$ (that is, if $(1-\varepsilon)k\leq 1$), then we can use $z:=x$.
+Thus, we may assume that $s<1$.
+For $f\in L(F(\operatorname{Cu}(A)))$, let $\alpha(f)\in\operatorname{Cu}(A)$ be as in Theorem 2.4, hence $\widehat{\alpha(f)}=f$.
+Set
+$z:=x\wedge\alpha\left(s\widehat{x}\right).$
+Then $z\leq x$. Further, using that $s<1$, and using [APRT22, Theorem 6.12], we have
+$\widehat{z}=\widehat{x}\wedge s\widehat{x}=s\widehat{x}=\frac{(1-\varepsilon)^{-1}}{k}\widehat{x}$
+and thus $(1-\varepsilon)k\widehat{z}=\widehat{x}$.
+Finally, let us verify that $x^{\prime}\leq kz$. In fact we will see that $x\leq kz$. Since $\operatorname{Cu}(A)$ is inf-semilattice ordered, we have
+$kz=k\big{(}x\wedge\alpha(s\widehat{x})\big{)}=\bigwedge_{i=0}^{k}\big{(}ix+(k-i)\alpha(s\widehat{x})\big{)}.$
+Clearly $x\leq ix+(k-i)\alpha(s\widehat{x})$ when $i>0$.
+For $i=0$, first note that $ks-1>0$ and therefore $\widehat{x}\leq\infty(ks-1)\widehat{x}$, which allows us to apply the partial additivity of $\alpha$ stated in Theorem 2.4.
+Hence, we get
+$x\leq x+\alpha\big{(}(ks-1)\widehat{x}\big{)}=\alpha\big{(}\widehat{x}+(ks-1)\widehat{x}\big{)}=\alpha(ks\widehat{x})=k\alpha(s\widehat{x}),$
+and thus $x\leq kz$, as desired. ∎
+Proposition 4.11.
+Let $A$ be a simple, unital, non-elementary $C^{*}$-algebra with a unique normalized quasitrace. Then $\operatorname{Cu}(A)$ is functionally divisible.
+Proof.
+By assumption, there exists a unique functional $\lambda\colon\operatorname{Cu}(A)\to[0,\infty]$ satisfying $\lambda([1])=1$.
+Note that any two elements $x,y\in\operatorname{Cu}(A)$ satisfy $\widehat{x}\leq\widehat{y}$ if and only if $\lambda(x)\leq\lambda(y)$.
+We first verify condition (iii) in Proposition 4.5, which then implies that $\operatorname{Cu}(A)$ satisfies condition (L) in Definition 4.3.
+Let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, and let $\varepsilon\in(0,1)$.
+We need to find $y\in\operatorname{Cu}(A)$ such that
+$2y\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\lambda(x^{\prime})\leq 2\lambda(y).$
+If $x^{\prime}=0$, then we can use $y=0$.
+Therefore, we may from now on assume that $x^{\prime}\neq 0$.
+Then $\lambda(x^{\prime})\in(0,\infty)$.
+Since $A$ is non-elementary, we can pick $v\in\operatorname{Cu}(A)$ such that $\lambda(v)=\tfrac{1}{2}\lambda(x^{\prime})$.
+We have
+$(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\frac{1}{2}\lambda(x^{\prime})\leq\lambda(x),\lambda(v).$
+Using that $\operatorname{Cu}(A)$ satisfies Edwards’ condition for $\lambda$ (see Theorem 4.7 and Remark 4.2(3) in [Thi20], and [APRT21]), we obtain $w\in\operatorname{Cu}(A)$ such that
+$(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\lambda(w),\,\,\,\text{ and }\,\,\,w\leq v,x.$
+Choose $w^{\prime\prime},w^{\prime}\in\operatorname{Cu}(A)$ such that
+$(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\lambda(w^{\prime\prime}),\,\,\,\text{ and }\,\,\,w^{\prime\prime}\ll w^{\prime}\ll w.$
+Applying (O5) for $w^{\prime\prime}\ll w^{\prime}\leq x$, we obtain $c\in\operatorname{Cu}(A)$ such that
+$w^{\prime\prime}+c\leq x\leq w^{\prime}+c.$
+Using that $\lambda(x^{\prime})\leq\lambda(x)$ and $\lambda(w^{\prime})\leq\lambda(w)\leq\lambda(v)=\tfrac{1}{2}\lambda(x^{\prime})$, we get
+$\frac{1}{2}\lambda(x^{\prime})=\lambda(x^{\prime})-\frac{1}{2}\lambda(x^{\prime})\leq\lambda(x)-\lambda(w^{\prime})\leq\lambda(c).$
+Then
+$(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\lambda(w^{\prime\prime}),\,\,\,\text{ and }\,\,\,(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\frac{1}{2}\lambda(x^{\prime})\leq\lambda(c).$
+Applying Edwards’ condition again, we obtain $y\in\operatorname{Cu}(A)$ such that
+$(1-\varepsilon)\frac{1}{2}\lambda(x^{\prime})<\lambda(y),\,\,\,\text{ and }\,\,\,y\leq w^{\prime\prime},c.$
+Then $2y\leq w^{\prime\prime}+c\leq x$. Thus $y$ has the desired properties.
+Next, to verify condition (U) of Definition 4.3, let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, let $k\in{\mathbb{N}}$ with $k\geq 1$, and let $\varepsilon\in(0,1)$.
+We need to find $z\in\operatorname{Cu}(A)$ such that
+$z\leq x,\quad(1-\varepsilon)k\lambda(z)\leq\lambda(x),\,\,\,\text{ and }\,\,\,x^{\prime}\leq kz.$
+Choose $\delta>0$ such that
+$1+(2k-1)\delta\leq(1-\varepsilon)^{-1}.$
+Then pick $v^{\prime\prime},v^{\prime},v\in\operatorname{Cu}(A)$ such that
+$x^{\prime}\ll v^{\prime\prime}\ll v^{\prime}\ll v\ll x,\,\,\,\text{ and }\,\,\,\lambda(v)\leq(1+\delta)\lambda(v^{\prime\prime}).$
+We have already verified condition (L) of Definition 4.3. Applied for $v^{\prime}\ll v$, for $k$ and for $\tfrac{\delta}{2}$, we obtain $y\in\operatorname{Cu}(A)$ such that
+$ky\leq v,\,\,\,\text{ and }\,\,\,\left(1-\frac{\delta}{2}\right)\lambda(v^{\prime})\leq k\lambda(y).$
+By Proposition 2.3(i), we have $(1-\delta)\widehat{v^{\prime\prime}}\ll(1-\tfrac{\delta}{2})\widehat{v^{\prime}}$ in $\operatorname{Lsc}(F(\operatorname{Cu}(A)))$, which allows us to choose $y^{\prime}\in\operatorname{Cu}(A)$ such that
+$y^{\prime}\ll y,\,\,\,\text{ and }\,\,\,(1-\delta)\lambda(v^{\prime\prime})\leq k\lambda(y^{\prime}).$
+Applying Lemma 2.1 for $ky\leq v$ and $y^{\prime}\ll y$, we obtain $z\in\operatorname{Cu}(A)$ such that
+$(k-1)y^{\prime}+z\leq v\leq(k-1)y+z,\,\,\,\text{ and }\,\,\,v\leq kz.$
+Using that $\lambda(y^{\prime})$ is finite, we get
+$k\lambda(z)$ $\leq k\lambda(v)-(k-1)k\lambda(y^{\prime})$
+$\leq k(1+\delta)\lambda(v^{\prime\prime})-(k-1)(1-\delta)\lambda(v^{\prime\prime})$
+$=(1+(2k-1)\delta)\lambda(v^{\prime\prime})\leq(1-\varepsilon)^{-1}\lambda(v^{\prime\prime}),$
+and thus
+$(1-\varepsilon)k\lambda(z)\leq\lambda(v^{\prime\prime})\leq\lambda(x).$
+We further have
+$z\leq v\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq v\leq kz,$
+which shows that $z$ has the claimed properties. ∎
 5. Pure C*-algebras
-In this section, we show that every $(m,n)$-pure $C^{*}$-algebra is pure; see Theorem 5.8.
+This section is devoted to proving Theorems A and C.
+We actually show that a $C^{*}$-algebra is pure if, and only if, it has controlled comparison and is functionally divisible; if, and only if, it is $(m,n)$-pure for some $m$ and $n$.
+Building on the results from Section 4 concerning automatic functional divisibility, we show that a $C^{*}$-algebra with controlled comparison is pure if it is either nowhere scattered with real rank zero or stable rank one, or also if it is simple, unital, non-elementary with a unique normalized quasitrace.
 Definition 5.1.
-Given $m,n\in{\mathbb{N}}$, we say that a $\mathrm{Cu}$-semigroup $S$ is $(m,n)$-pure if $S$ has $m$-comparison (Definition 3.1) and is $n$-almost divisible (Definition 4.1).
-We say that $S$ is pure if it is $(0,0)$-pure.
-A $\mathrm{Cu}$-semigroup $S$ has strict comparison if, whenever $x,y\in S$ with $x\leq\infty y$ satisfy that $\lambda(x)<\lambda(y)$ for any $\lambda\in F(S)$ normalized at $y$, one has $x\leq y$.
-A $C^{*}$-algebra is $(m,n)$-pure if its Cuntz semigroup is.
-Similarly, a $C^{*}$-algebra is pure if it is $(0,0)$-pure.
-We also say that a $C^{*}$-algebra has strict comparison of positive elements (by $2$-quasitraces) provided its Cuntz semigroup has strict comparison.
-It was proved in [ERS11, Proposition 6.2] that $S$ has strict comparison if and only if it is almost unperforated.
+Given $m,n\in{\mathbb{N}}$, we say that a $\mathrm{Cu}$-semigroup is $(m,n)$-pure if it has $m$-comparison and is $n$-almost divisible.
+We say that a $\mathrm{Cu}$-semigroup is pure if it is $(0,0)$-pure.
+A $C^{*}$-algebra is $(m,n)$-pure if its Cuntz semigroup is. Similarly, a $C^{*}$-algebra is pure if it is $(0,0)$-pure.
 Proposition 5.2.
-Every $\mathcal{Z}$-stable $C^{*}$-algebra is pure. Every pure $C^{*}$-algebra has strict comparison of positive elements.
+Every $\mathcal{Z}$-stable $C^{*}$-algebra is pure. Every pure $C^{*}$-algebra has the Global Glimm Property and strict comparison (of positive elements by quasitraces).
 Proof.
 Let $A$ be a $\mathcal{Z}$-stable $C^{*}$-algebra.
-It was shown in [Rør04, Theorem 4.5] that $\operatorname{Cu}(A)$ is almost unperforated, and thus by the comment above $A$ has strict comparison of positive elements.
-The argument used to prove almost unperforation can be used to show almost divisibility.
-For a proof, see e.g.
-[APT11, Theorem 5.35].
+Then the (classical, uncomplete) Cuntz semigroup $W(A)$ is almost unperforated by [Rør04, Theorem 4.5].
+Further, by [APT11, Theorem 5.35], for every $x\in W(A)$ and every $k\geq 1$ there exists $y\in W(A)$ such that $ky\leq x\leq(k+1)y$.
+Using that $W(A)$ is order-dense in $\operatorname{Cu}(A)$ (see [APT18, Theorem 3.2.8]), it follows that $\operatorname{Cu}(A)$ is almost unperforated and almost divisible (see also Paragraph 4.2), and thus pure.
+By definition, every pure $C^{*}$-algebra is almost divisible, and therefore $(2,\omega)$-divisible, which by [TV23, Theorem 3.6] is equivalent to the Global Glimm Property.
+Further, every pure $C^{*}$-algebra has 00-comparison, and thus enjoys strict comparison as noted in Paragraph 3.4.
 ∎
-Remark 5.3.
-It is perhaps worth pointing out that the results in [Rør04] deal with the uncomplete Cuntz semigroup $\mathrm{W}(A)$.
-This makes no difference regarding almost unperforation, since $\operatorname{Cu}(A)$ is a completion of $\mathrm{W}(A)$ (see [APT18, Theorem 3.2.8]), hence $\mathrm{W}(A)$ is almost unperforated if and only if $\operatorname{Cu}(A)$ is.
 One can ask whether, for $\mathrm{Cu}$-semigroups, $(0,0)$-pureness agrees with some sort of tensorial absorption.
@@
 (See also [APP18, Section 7].) The following question is thus pertinent:
-Question 5.5.
+Question 5.4.
 Let $A$ be a pure $C^{*}$-algebra. Is $\operatorname{Cu}(A)\cong\operatorname{Cu}(\mathcal{Z}\otimes A)$?
@@
 Proposition 5.5.
-Let $(S,\Sigma)$ be a scaled $\mathrm{Cu}$-semigroup satisfying (O5) and (O7).
-Assume that $S$ has locally bounded comparison amplitude, and also that is upper and lower functionally almost divisible.
-Then $S$ is pure.
+Let $(S,\Sigma)$ be a scaled $\mathrm{Cu}$-semigroup satisfying (O5). Assume that $(S,\Sigma)$ has controlled comparison and is functionally divisible. Then $S$ is pure.
 Proof.
@@
 This is equivalent to showing that $x^{\prime}\leq y$ for any $x^{\prime}\in S$ such that $x^{\prime}\ll x$.
-Therefore, let $x^{\prime}\ll x$ and pick $x^{\prime\prime}\in S$ such that $x^{\prime}\ll x^{\prime\prime}\ll x$.
+Therefore, let $x^{\prime}\in S$ with $x^{\prime}\ll x$, and pick $x^{\prime\prime}\in S$ such that $x^{\prime}\ll x^{\prime\prime}\ll x$.
 Since we have $\widehat{x}\leq\tfrac{n}{n+1}\widehat{y}$, we may choose $\gamma,\gamma^{\prime}$ such that $\tfrac{n}{n+1}<\gamma<\gamma^{\prime}<1$.
@@
 Take $y^{\prime\prime}$ such that $y^{\prime}\ll y^{\prime\prime}\ll y$, and choose $d\in{\mathbb{N}}$ such that $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$.
-Since $(S,\Sigma)$ has locally bounded comparison amplitude, we obtain $N=N(\gamma^{\prime},d)\in{\mathbb{N}}$ such that
-$\widehat{v}\leq\gamma^{\prime}\widehat{w}\quad\Rightarrow\quad v\leq_{N}w,\quad\text{ for all }v,w\in\Sigma^{(d)}.$
-Let $M\in{\mathbb{N}}$ be such that $M\geq 2$, $M\geq N$ and $\gamma\tfrac{M+1}{M-1}\leq\gamma^{\prime}$. Using (UFAD) for $x^{\prime}\ll x^{\prime\prime}$ and $M-1$, we obtain $v\in S$ satisfying
-$v\leq x^{\prime\prime},\quad(M-1)\widehat{v}\leq\widehat{x^{\prime\prime}},\,\,\,\text{ and }\,\,\,x^{\prime}\leq Mv.$
-Applying (LFAD) for $y^{\prime}\ll y^{\prime\prime}$ and $M$, one gets $w\in S$ such that
-$Mw\leq y^{\prime\prime},\,\,\,\text{ and }\,\,\,\widehat{y^{\prime}}\leq(M+1)\widehat{w}.$
+Since $(S,\Sigma)$ has controlled comparison, we obtain $N=N(\gamma^{\prime},d)\in{\mathbb{N}}$ such that
+$\widehat{v}\leq\gamma^{\prime}\widehat{w}\quad\text{ implies }\quad v\leq_{N}w,\quad\text{ for all }v,w\in\Sigma^{(d)}.$
+Pick $\varepsilon>0$ such that $\gamma\leq(1-\varepsilon)^{2}\gamma^{\prime}$. Applying condition (U) in Definition 4.3 for $x^{\prime}\ll x^{\prime\prime}$ and $N$, we obtain $v\in S$ satisfying
+$v\leq x^{\prime\prime},\quad(1-\varepsilon)N\widehat{v}\leq\widehat{x^{\prime\prime}},\,\,\,\text{ and }\,\,\,x^{\prime}\leq Nv.$
+Similarly, applying condition (L) in Definition 4.3 for $y^{\prime}\ll y^{\prime\prime}$ and $N$, one gets $w\in S$ such that
+$Nw\leq y^{\prime\prime},\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{y^{\prime}}\leq N\widehat{w}.$
 Combining the functional inequalities for $\widehat{v},\widehat{w}$ and using that $\widehat{x^{\prime\prime}}\leq\gamma\widehat{y^{\prime}}$ at the second step, we have
-$(M-1)\widehat{v}\leq\widehat{x^{\prime\prime}}\leq\gamma\widehat{y^{\prime}}\leq\gamma(M+1)\widehat{w},$
-which implies, by our choice of $M$, that $\widehat{v}\leq\gamma^{\prime}\widehat{w}$.
-Note that both $v$ and $w$ belong to $\Sigma^{(d)}$, since $v\leq x^{\prime\prime},Mw\leq y^{\prime\prime}$, and $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$.
-We deduce that $v\leq_{N}w$ and, since $M\geq N$, we get $Mv\leq Mw$.
-Hence,
-$x^{\prime}\leq Mv\leq Mw\leq y^{\prime\prime}\leq y,$
+$(1-\varepsilon)^{2}N\widehat{v}\leq(1-\varepsilon)\widehat{x^{\prime\prime}}\leq(1-\varepsilon)\gamma\widehat{y^{\prime}}\leq\gamma N\widehat{w}\leq(1-\varepsilon)^{2}N\gamma^{\prime}\widehat{w},$
+which implies that $\widehat{v}\leq\gamma^{\prime}\widehat{w}$.
+Note that both $v$ and $w$ belong to $\Sigma^{(d)}$, since $v\leq x^{\prime\prime},Nw\leq y^{\prime\prime}$, and $x^{\prime\prime},y^{\prime\prime}\in\Sigma^{(d)}$.
+It follows that $v\leq_{N}w$, and thus
+$x^{\prime}\leq Nv\leq Nw\leq y^{\prime\prime}\leq y,$
 as desired.
-To show that $S$ is almost divisible, let $k\geq 2$ and take $x^{\prime},x\in S$ such that $x^{\prime}\ll x$. Since $S$ satisfies (LFAD), we may apply Lemma 4.4(ii) to obtain $y\in S$ such that
-$ky\leq x,\,\,\,\text{ and }\,\,\,\widehat{x^{\prime}}\leq\left(k+\frac{1}{k}\right)\widehat{y}.$
+To show that $S$ is almost divisible, let $x^{\prime}\ll x$ in $S$, and let $k\in{\mathbb{N}}$.
+Pick $\varepsilon\in(0,1)$ such that $(1-\varepsilon)^{-1}k<k+1$.
+Applying condition (L) in Definition 4.3 for $x^{\prime\prime}\ll x$, and for $k$ and $\varepsilon$, we obtain $y\in S$ such that
+$ky\leq x,\,\,\,\text{ and }\,\,\,(1-\varepsilon)\widehat{x^{\prime}}\leq k\widehat{y}.$
 It follows that $x^{\prime}\leq\infty y$; see, for example, [APRT22, Lemma 6.6(iii)].
-We also have that $\lambda(x^{\prime})<\lambda((k+1)y)$ for every functional normalized at $y$.
+Since $(1-\varepsilon)^{-1}k<k+1$, we also have that $\lambda(x^{\prime})<\lambda((k+1)y)$ for every functional normalized at $y$.
 Since $S$ is almost unperforated, we obtain $x^{\prime}\leq(k+1)y$ by [ERS11, Proposition 6.2].
 ∎
+Remark 5.6.
+It is possible to adapt the proof of Proposition 5.5 so that instead of (U) it suffices to assume the condition ($U_{2^{m}}$) from Remark 4.6 for all $m$.
+Namely, the following stronger form of Proposition 5.5 is available: If a scaled $\mathrm{Cu}$-semigroup $(S,\Sigma)$ satisfies (O5) and has controlled comparison, and for every $x^{\prime}\ll x$ in $S$ and every $\varepsilon\in(0,1)$ there exist $y,z\in S$ such that
+$2y\leq x,\quad(1-\varepsilon)\widehat{x^{\prime}}\leq 2\widehat{y},\quad z\leq x,\quad(1-\varepsilon)2\widehat{z}\leq\widehat{x},\,\,\,\text{ and }\,\,\,x^{\prime}\leq 2z,$
+then $S$ is pure.
 Theorem 5.7.
-Let $A$ be a $C^{*}$-algebra that has locally bounded comparison amplitude and is $n$-almost divisible for some $n\in{\mathbb{N}}$. Then $A$ is pure.
+Given a $C^{*}$-algebra $A$, the following statements are equivalent:
+  1. (i)
+    $A$ is pure, that is, $A$ has strict comparison and is almost divisible.
+  2. (ii)
+    $A$ is $(m,n)$-pure for some $m,n\in{\mathbb{N}}$.
+  3. (iii)
+    $A$ has controlled comparison and is functionally divisible.
 Proof.
-Since $\operatorname{Cu}(A)$ is $n$-almost divisible, Theorem 4.7 gives that $\operatorname{Cu}(A)$ has (LFAD) and (UFAD), and the result follows from Proposition 5.6. ∎
+It is clear that (i) implies (ii).
+Further, (ii) implies (iii) since $m$-comparison implies controlled comparison by Proposition 3.7, and $n$-almost divisibility implies functional divisibility by Proposition 4.9.
+Finally, (iii) implies (i) by Proposition 5.5 since the Cuntz semigroup of every $C^{*}$-algebra satisfies (O5).
+∎
+Remark 5.8.
+If follows from Theorem 5.7 (see also Paragraph 4.2) that $A$ is $(m,n)$-pure if, and only if, $\operatorname{Cu}(A)$ has $m$-comparison and is $n$-almost divisible in the sense of [Win12, Definition 3.5], that is, given $x\in\operatorname{Cu}(A)$ and $k\geq 1$ , there is $y\in\operatorname{Cu}(A)$ such that $ky\leq x\leq(n+1)(k+1)y$.
+Remark 5.9.
+Theorem 5.7 shows that the combination of weak forms of comparison and divisibility implies strong forms of comparison and divisibility.
+One may wonder if this already holds individually for comparison and divisibility.
+For the comparison properties, this is not the case.
+Indeed, there exist $C^{*}$-algebras that have controlled comparison (or even $m$-comparison for some $m\geq 1$) that do not have 00-comparison.
+For example, if $X=[0,1]^{5}$ is the five-dimensional cube, then $C(X)$ has nuclear dimension at most five, and therefore has $5$-comparison by [Rob11, Theorem 1.3], but it does not have 00-comparison since the radius of comparison is at least one by [EN13, Theorem 1.1].
+Similarly, we expect that the answer to the following question is positive.
+Question 5.10.
+Does there exist a $C^{*}$-algebra that is functionally divisible (or even $n$-almost divisible for some $n$) but not almost divisible?
+Proposition 5.11.
+Let $A$ be a functionally divisible $C^{*}$-algebra. Then the following statements are equivalent:
+  1. (i)
+    $A$ is pure.
+  2. (ii)
+    $A$ has strict comparison (of positive elements by quasitraces).
+  3. (iii)
+    $A$ has $m$-comparison for some $m$.
+  4. (iv)
+    $A$ has controlled comparison.
+Proof.
+By definition, pureness implies 00-comparison, which is equivalent to strict comparison by Paragraph 3.4.
+This shows that (i) implies (ii), which in turn implies (iii).
+Further, (iii) implies (iv) by Proposition 3.7.
+Finally, (iv) implies (i) by Theorem 5.7.
+∎
+Combining Proposition 5.11 with the results proving functional divisibility from Section 4, we can now verify pureness in various settings:
 Theorem 5.12.
-If a $C^{*}$-algebra is $(m,n)$-pure for some $m,n\in{\mathbb{N}}$, then it is pure.
+Let $A$ be a nowhere scattered $C^{*}$-algebra that has real rank zero or stable rank one. Assume that $A$ has controlled comparison (for example, $m$-comparison for some $m$). Then $A$ is pure.
 Proof.
-If $A$ is $(m,n)$-pure, then since it has $m$-comparison Lemma 3.6 implies that $A$ has locally bounded comparison amplitude.
-Since it is also $n$-almost divisible, we may apply Theorem 5.7 to conclude that $A$ is pure.
+If $A$ is nowhere scattered and has real rank zero, then $A$ is almost divisible as noted in Paragraph 4.2.
+If $A$ is nowhere scattered and has stable rank one, then $A$ is functionally divisible by Proposition 4.10.
+Now, in both cases the result follows from Proposition 5.11.
 ∎
+Theorem 5.13.
+Let $A$ be a simple, unital, non-elementary $C^{*}$-algebra with a unique normalized quasitrace. Assume that $A$ has controlled comparison (for example, $m$-comparison for some $m$). Then $A$ is pure.
+Proof.
+By Proposition 4.11, $\operatorname{Cu}(A)$ is functionally divisible. Now, the result follows from Proposition 5.11. ∎
+Example 5.14.
+Let $G$ be an infinite, discrete group such that the reduced group $C^{*}$-algebra $C^{*}_{\mathrm{red}}(G)$ is simple.
+Then $C^{*}_{\mathrm{red}}(G)$ is simple, unital, non-elementary and has a unique trace by [BKKO17, Corollary 4.3].
+If $G$ is exact, then quasitraces on $C^{*}_{\mathrm{red}}(G)$ are traces by Haagerup’s theorem ([Haa14]), and it follows that $C^{*}_{\mathrm{red}}(G)$ has a unique normalized quasitrace.
+In this case, Theorem 5.13 applies, and it follows that $C^{*}_{\mathrm{red}}(G)$ is pure if, and only if, it has controlled comparison.
+It has been proved very recently in [AGKP24] that, for $\mathbb{F}_{n}$, the free group on $n$ generators, the reduced group $C^{*}$-algebra $C^{*}_{\mathrm{red}}(\mathbb{F}_{n})$ has strict comparison (equivalently, if it is pure).
+The result in fact holds for a large class of discrete groups.
+For non-exact groups, we ask the following question:
+Question 5.15.
+Do simple reduced group $C^{*}$-algebras have a unique normalized quasitrace?
+Example 5.16.
+In his groundbreaking work on regularity properties for simple, nuclear $C^{*}$-algebras, Villadsen constructed examples of such algebras that fail strict comparison, and consequently are not pure.
+The algebras of ‘first type’ from [Vil98] have stable rank one (but possibly a complicated trace simplex), while the algebras of ‘second type’ from [Vil99] have higher stable rank but a unique tracial state (and consequently also a unique quasitracial state).
+Applying Propositions 4.10 and 4.11, we see that both types of Villadsen algebras are automatically functionally divisible.
+We deduce that Villadsen algebras that fail strict comparison also fail the much weaker property of controlled comparison (and in particular, they do not have $m$-comparison for any $m$).
 6. C*-algebras with the Global Glimm Property and finite nuclear dimension
-The next result is our variation of [RT17, Proposition 3.2(ii)] for nowhere scattered $C^{*}$-algebras. This has essentially appeared in the argument of [Vil23, Proposition 4.1].
+In this section, we prove Theorem B.
+For this, we use that a $C^{*}$-algebra has the Global Glimm Property precisely when its Cuntz semigroup is $(2,\omega)$-divisible; see [TV23, Theorem 3.6].
+We then verify a particular statement of the non-simple Toms-Winter conjecture for separable, locally subhomogeneous $C^{*}$-algebras with stable rank one and topological dimension zero, which yields Theorem E.
+The next result is a variation of [RT17, Proposition 3.2(ii)] for nowhere scattered $C^{*}$-algebras.
+This has essentially appeared in the argument of [Vil23, Proposition 4.1].
+We use $\dim_{\mathrm{nuc}}(A)$ to denote the nuclear dimension of $A$.
 Lemma 6.1.
-Let $A$ be a nowhere scattered $C^{*}$-algebra with $\dim_{\mathrm{nuc}}(A)\leq m$.
-Then for every $k\geq 1$, and every pair $x^{\prime},x\in\operatorname{Cu}(A)$ satisfying $x^{\prime}\ll x$, there exists $y\in\operatorname{Cu}(A)$ such that
+Let $m\in{\mathbb{N}}$, and let $A$ be a nowhere scattered $C^{*}$-algebra satisfying $\dim_{\mathrm{nuc}}(A)\leq m$.
+Then for every $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$ and every $k\in{\mathbb{N}}$ with $k\geq 1$, there exists $y\in\operatorname{Cu}(A)$ such that
 $y\ll x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll ky\ll 2(m+1)x.$
 Proof.
-By [TV24, Proposition 4.12] and [WZ10, Proposition 2.3], $A\otimes\mathcal{K}$ is nowhere scattered and of nuclear dimension at most $m$. We may thus assume that $A$ is stable.
+By [TV24a, Proposition 4.12] and [WZ10, Proposition 2.3], $A\otimes\mathcal{K}$ is nowhere scattered and of nuclear dimension at most $m$. We may thus assume that $A$ is stable.
 Choose $a\in A_{+}$ and $\varepsilon>0$ such that
 $x^{\prime}\ll[(a-\varepsilon)_{+}],\,\,\,\text{ and }\,\,\,[a]\ll x.$
-Since $A$ is nowhere scattered, the hereditary sub-$C^{*}$-algebra $\overline{aAa}$ has no finite-dimensional representations, by [TV24, Theorem 3.1].
+Since $A$ is nowhere scattered, the hereditary sub-$C^{*}$-algebra $\overline{aAa}$ has no finite-dimensional representations, by [TV24a, Theorem 3.1].
 It also has nuclear dimension at most $m$ (by [WZ10, Proposition 2.5]), whence we can apply [RT17, Proposition 3.2(ii)] for $k$ (and for $\overline{aAa}$, $a$ and $\varepsilon$) to obtain $b\in\overline{aAa}_{+}$ such that
@@
 Then $y:=[b]\in\operatorname{Cu}(A)$ has the desired properties. ∎
-The following is inspired by [RT17, Lemma 3.4].
+The next two results are inspired by [RT17, Lemma 3.4].
 Lemma 6.2.
-Let $A$ be a $C^{*}$-algebra with the Global Glimm Property, and let $m,l\in{\mathbb{N}}$.
-Assume that $\dim_{\mathrm{nuc}}A\leq m$.
-Then, there exists $L=L(m,l)$ such that, for every $x^{\prime}\ll x$, there exist $y_{0},\ldots,y_{l}\in\operatorname{Cu}(A)$ such that
-$y_{0}+\ldots+y_{l}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll Ly_{j}$
-for $j=0,\ldots,l$.
+Let $m\in{\mathbb{N}}$, and let $A$ be a $C^{*}$-algebra with the Global Glimm Property and with $\dim_{\mathrm{nuc}}(A)\leq m$.
+Set $L:=14m^{2}+6$.
+Then for every $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, there exist $y_{0},y_{1}\in\operatorname{Cu}(A)$ such that
+$y_{0}+y_{1}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll Ly_{0},Ly_{1}.$
 Proof.
-Let $x^{\prime},x\in\operatorname{Cu}(A)$ be such that $x^{\prime}\ll x$. We claim that there exists a separable sub-$C^{*}$-algebra $B\subseteq A$ such that
-  • (i)
-    $B$ has the Global Glimm Property;
-  • (ii)
-    $\dim_{\mathrm{nuc}}(B)\leq m$;
-  • (iii)
+Let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$. We claim that there exists a separable sub-$C^{*}$-algebra $B\subseteq A$ such that
+  1. (i)
+    $B$ has the Global Glimm Property and $\dim_{\mathrm{nuc}}(B)\leq m$;
+  2. (ii)
     the induced inclusion $\operatorname{Cu}(\iota)\colon\operatorname{Cu}(B)\to\operatorname{Cu}(A)$ is an order-embedding;
-  • (iv)
+  3. (iii)
     $x^{\prime}$ and $x$ are in the image of $\operatorname{Cu}(\iota)$.
@@
 Now [TV21, Proposition 6.1] allows us to choose a sub-$C^{*}$-algebra $B$ with the required conditions.
-Therefore we may assume that $A$ is separable and that, as in the proof of Lemma 6.1 above, we may also assume that $A$ is stable.
-Now we apply induction on $l$. Note that, if $l=0$, one can take $y_{0}:=x$. Let us assume that $l=1$.
+Identifying $\operatorname{Cu}(B)$ with its image in $\operatorname{Cu}(A)$, we view $x^{\prime},x$ as elements in $\operatorname{Cu}(B)$, and it suffices to find $y_{0},y_{1}\in\operatorname{Cu}(B)$ with $y_{0}+y_{1}\leq x$ and $x^{\prime}\ll Ly_{0},Ly_{1}$.
+Therefore we may assume that $A$ is separable.
 Choose $x^{\prime\prime}\in\operatorname{Cu}(A)$ such that $x^{\prime}\ll x^{\prime\prime}\ll x$.
@@
 Further,
-$x^{\prime}\ll 2e^{\prime\prime}\ll 2(2m+3)f^{\prime}=(4m+6)y_{0}.$
+$x^{\prime}\ll 2e^{\prime\prime}\ll 2(2m+3)f^{\prime}=(4m+6)y_{0}\leq(14m^{2}+6)y_{0}=Ly_{0}.$
 Multiplying (2) by $32m+$ 3 at the first step, and using (1) at the second step, we deduce
@@
 $\lambda(e^{\prime})\leq\infty=(2m+3)\lambda(c^{\prime})\leq(2m+3)\lambda(y_{1}).$
-Since $\operatorname{Cu}(A)$ has $m$-comparison (see Remarks 3.2 (1)), we get
-$e^{\prime}\leq 2m(2m+3)y_{1}.$
+Since $\operatorname{Cu}(A)$ has $m$-comparison (see Remarks 3.2), we get
+$e^{\prime}\leq(m+1)(2m+3)y_{1}.$
 Using that $x^{\prime}\ll 2e^{\prime}$, one has
-$x^{\prime}\ll 2e^{\prime}\leq 4m(2m+3)y_{1}.$
-Thus, setting $L:=\max\{4m+6,4m(2m+3)\}$, we see that $y_{0}+y_{1}\leq x$ and $x^{\prime}\ll Ly_{0},Ly_{1}$. This shows the statement for $l=1$.
-Now let $l\in{\mathbb{N}}$ and assume that we have proven the statement for $l-1$.
-Take $x^{\prime},x\in\operatorname{Cu}(A)$ such that $x^{\prime}\ll x$. We get $L^{\prime}\in{\mathbb{N}}$ and elements $z_{0},\ldots,z_{l-1}$ such that
-$z_{0}+\ldots+z_{l-1}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll L^{\prime}z_{j}$
+$x^{\prime}\ll 2e^{\prime}\leq 2(m+1)(2m+3)y_{1}=2(2m^{2}+5m+3)y_{1}\leq(14m^{2}+6)y_{1}=Ly_{1}.$
+This shows that $y_{0}$ and $y_{1}$ have the desired properties. ∎
+Lemma 6.3.
+Given $m,l\in{\mathbb{N}}$, set $L:=(14m^{2}+6)^{l}$.
+Then, for every $C^{*}$-algebra $A$ with the Global Glimm Property and with $\dim_{\mathrm{nuc}}(A)\leq m$, and for every $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, there exist $y_{0},\ldots,y_{l}\in\operatorname{Cu}(A)$ such that
+$y_{0}+\ldots+y_{l}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll Ly_{j}$
+for $j=0,\ldots,l$.
+Proof.
+Fix some $m\in{\mathbb{N}}$, and let $A$ be a $C^{*}$-algebra that has the Global Glimm Property and with $\dim_{\mathrm{nuc}}(A)\leq m$. Write $L(m,l):=(14m^{2}+6)^{l}$ for $l\geq 0$.
+By induction over $l$, we verify that the following statement holds:
+  1. ($D_{l}$)
+    For every $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$ there exist $y_{0},\ldots,y_{l}\in\operatorname{Cu}(A)$ such that
+    $y_{0}+\ldots+y_{l}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll L(m,l)y_{j}$
+    for $j=0,\ldots,l$.
+To verify ($D_{0}$), given $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$ we simply use $y_{0}=x$.
+Further, ($D_{1}$) was shown in Lemma 6.2.
+Now let $l\in{\mathbb{N}}$ with $l\geq 2$, and assume that we have proved ($D_{l-1}$).
+To verify ($D_{l}$), let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$.
+Applying ($D_{l-1}$) for $x^{\prime}\ll x$, we get $z_{0},\ldots,z_{l-1}\in\operatorname{Cu}(A)$ such that
+$z_{0}+\ldots+z_{l-1}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\ll L(m,l-1)z_{j}$
 for $j=0,\ldots,l-1$.
-Set $y_{j}:=z_{j}$ for each $j\leq l-2$, and take $z_{l-1}^{\prime}\in\operatorname{Cu}(A)$ such that $z_{l-1}^{\prime}\ll z_{l-1}$ and $x^{\prime}\ll L^{\prime}z_{l-1}^{\prime}$.
-By the argument above, one finds $L^{\prime\prime}\in{\mathbb{N}}$ and $y_{l-1},y_{l}\in\operatorname{Cu}(A)$ satisfying
-$y_{l-1}+y_{l}\leq z_{l-1},\,\,\,\text{ and }\,\,\,z_{l-1}^{\prime}\ll L^{\prime\prime}y_{l-1},L^{\prime\prime}y_{l}.$
-One can now check that the elements $y_{0},\ldots,y_{l}$ and $L:=L^{\prime}L^{\prime\prime}$ satisfy the required conditions. ∎
+Set $y_{j}:=z_{j}$ for $j=0,\ldots,l-2$, and take $z_{l-1}^{\prime}\in\operatorname{Cu}(A)$ such that
+$z_{l-1}^{\prime}\ll z_{l-1},\,\,\,\text{ and }\,\,\,x^{\prime}\ll L(m,l-1)z_{l-1}^{\prime}.$
+Applying ($D_{1}$) for $z_{l-1}^{\prime}\ll z_{l-1}$, we find $y_{l-1},y_{l}\in\operatorname{Cu}(A)$ satisfying
+$y_{l-1}+y_{l}\leq z_{l-1},\,\,\,\text{ and }\,\,\,z_{l-1}^{\prime}\ll L(m,1)y_{l-1},L(m,1)y_{l}.$
+Then $y_{0}+\ldots+y_{l}\leq x$, and
+$x^{\prime}\ll L(m,l-1)z_{j}\leq L(m,l)z_{j}=L(m,l)y_{j}$
+for $j=0,\ldots,l-2$ and
+$x^{\prime}\ll L(m,l-1)z_{l-1}^{\prime}\ll L(m,l-1)L(m,1)y_{j}=L(m,l)y_{j}$
+for $j=l-1,l$. This shows that $y_{0},\ldots,y_{l}$ have the desired properties. ∎
 Proposition 6.4.
-Let $A$ be a $C^{*}$-algebra with the Global Glimm Property, and let $m\in{\mathbb{N}}$.
-Assume that $\dim_{\mathrm{nuc}}A\leq m$.
-Then, there exists $N=N(m)$ such that $\operatorname{Cu}(A)$ is $N$-almost divisible.
+Given $m\in{\mathbb{N}}$, there exists $N=N(m)$ with the following property: For every $C^{*}$-algebra $A$ with the Global Glimm Property and $\dim_{\mathrm{nuc}}(A)\leq m$, the Cuntz semigroup $\operatorname{Cu}(A)$ is $N$-almost divisible.
 Proof.
-The proof is inspired by that of [RT17, Lemma 3.6]. Let $L$ be the constant obtained from Lemma 6.2 for $m$ and $l=m$. We verify the statement for $N:=3L(m+1)$.
-Let $x^{\prime}\ll x$ and $k\geq 1$. We need to find $y$ such that
+The proof is inspired by that of [RT17, Lemma 3.6]. Let $L=L(m,m)$ be the constant obtained from Lemma 6.3. We verify the statement for $N:=3L(m+1)$.
+Let $A$ be a $C^{*}$-algebra with the Global Glimm Property and $\dim_{\mathrm{nuc}}(A)\leq m$.
+To show that $\operatorname{Cu}(A)$ is $N$-almost divisible, let $x^{\prime}\ll x$ in $\operatorname{Cu}(A)$, and let $k\geq 1$.
+We need to find $y\in\operatorname{Cu}(A)$ such that
 $ky\leq x,\,\,\,\text{ and }\,\,\,x^{\prime}\leq(k+1)(N+1)y.$
@@
 Applying Lemma 6.3 for $x^{\prime\prime}\ll x$, we obtain $y_{0},\ldots,y_{m}$ such that
-$y_{0}+\ldots+y_{m}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime\prime}\leq Ly_{j}\quad\text{ for $j=0,\ldots,m$}.$
+$y_{0}+\ldots+y_{m}\leq x,\,\,\,\text{ and }\,\,\,x^{\prime\prime}\ll Ly_{j}\quad\text{ for $j=0,\ldots,m$}.$
 For each $j$, we obtain
 $3kL(m+1)c^{\prime}\ll 2(m+1)x^{\prime\prime}\ll 2L(m+1)y_{j}.$
-and thus $k\widehat{c^{\prime}}<_{s}\widehat{y_{j}}$.
-By $m$-comparison, we get
+and thus $kc^{\prime}<_{s}y_{j}$.
+As noted in Remarks 3.2, $A$ has $m$-comparison, and we get
 $kc^{\prime}\leq y_{0}+\ldots+y_{m}\leq x.$
@@
 Proof.
-Let $A$ be a $C^{*}$-algebra with the Global Glimm Property and that has nuclear dimension at most $m$.
-By Remarks 3.2 (1), we have that $A$ has $m$-comparison.
-Using Proposition 6.3, we know that $A$ is $n$-almost divisible for some $n$.
+Let $m\in{\mathbb{N}}$, and let $A$ be a $C^{*}$-algebra with the Global Glimm Property and $\dim_{\mathrm{nuc}}(A)\leq m$.
+Then $A$ has $m$-comparison, as noted in Remarks 3.2.
+Further, by Proposition 6.4, we know that $A$ is $n$-almost divisible for some $n$.
 Hence, $A$ is $(m,n)$-pure, and the result now follows from Theorem 5.7.
 ∎
-Remark 6.5.
-In [RT17, Theorem 3.1] it is shown that if a nowhere scattered $C^{*}$-algebra $A$ has finite nuclear dimension and no quotient of $A$ has a simple, purely infinite ideal, then $A$ is pure.
-Further, every approximately subhomogeneous $C^{*}$-algebra (ASH-algebra) has locally finite nuclear dimension [NW06].
-By [ENST20, Theorem A], every $\mathcal{Z}$-stable ASH-algebra has decomposition rank at most two, and therefore finite nuclear dimension.
-Combining these known results, we obtain another partial positive answer to Question B: For a separable, nowhere scattered ASH-algebra whose primitive ideal space has a basis of compact-open subsets or is Hausdorff, the properties ‘finite nuclear dimension’, ‘$\mathcal{Z}$-stability’ and ‘pureness’ are equivalent.
-If one additionally assumes that the algebra has stable rank one, then these properties are also equivalent to ‘strict comparison of positive elements’.
+Theorem 6.6.
+Let $A$ be a separable, locally subhomogeneous $C^{*}$-algebra that has stable rank one and topological dimension zero. Then the following statements are equivalent:
+  1. (1)
+    $A$ has the Global Glimm Property and finite nuclear dimension.
+  2. (2)
+    $A$ is $\mathcal{Z}$-stable.
+  3. (3a)
+    $A$ is pure.
+  4. (3b)
+    $A$ is nowhere scattered and has strict comparison of positive elements.
+Proof.
+By Theorem 6.5, (1) implies (3a). By Proposition 5.2, (2) implies (3a), and (3a) implies (3b). By Theorem 5.12, (3b) implies (3a).
+Every locally subhomogeneous $C^{*}$-algebra has locally finite nuclear dimension by [NW06].
+Since locally subhomogeneous $C^{*}$-algebras have no simple, purely infinite ideal-quotients, it follows from [RT17, Theorem 7.10] that (3a) implies (2).
+Finally, by [ENST20, Theorem A], every $\mathcal{Z}$-stable, locally subhomogeneous $C^{*}$-algebra has decomposition rank at most two, and therefore finite nuclear dimension.
+Further, by Proposition 5.2, $\mathcal{Z}$-stability implies the Global Glimm Property.
+This shows that (2) implies (1).
+∎